### v2

# Classical facets of quantum integrability

Classical facets of quantum integrability
A. Zabrodin Skolkovo Institute of Science and Technology, 143026, Moscow, Russia and National Research University Higher School of Economics, 20 Myasnitskaya Ulitsa, Moscow 101000, Russia, and NRC “Kurchatov institute”, Moscow, Russia; e-mail: zabrodin@itep.ru
(January 2025)
ITEP-TH-06/25
This paper is a review of the works devoted to understanding and reinterpretation of the theory of quantum integrable models solvable by Bethe ansatz in terms of the theory of purely classical soliton equations. Remarkably, studying polynomial solutions of the latter by methods of classical soliton theory, one is able to develop a method of solving the spectral problem for the former which provides an alternative to the Bethe ansatz procedure. Our main examples are the generalized inhomogeneous spins chains with twisted boundary conditions on the quantum side and the modified Kadomtsev-Petviashvili hierarchy of nonlinear differential-difference equations on the classical side. In this paper, we restrict ourselves to quantum spin chains with rational $GL(n)$-invariant $R$-matrices (of the XXX type). Also, the connection of quantum spin chains with classical soliton equations implies a close interrelation between the spectral problem for spin chains and integrable many-body systems of classical mechanics such as Calogero-Moser and Ruijsenaars-Scheider models, which is known as the quantum-classical duality. Revisiting this topic, we suggest a simpler and more instructive proof of this kind of duality.
1 Introduction
Integrable models of different types, from mechanical systems with a finite number of degrees of freedom to models of field theory, play an outstanding role in modern mathematical physics. As a rule, they have important applications to physical problems as well as deep and beautiful mathematical structures and symmetries underlying their integrability. That is why integrable models are interesting from both physical and mathematical points of view.
An intriguing phenomenon in the world of integrable models, which already has been observed in many examples, is that the models belonging to very different classes (for example, such as mechanical systems, spin chains, nonlinear partial differential equations) have nontrivial hidden interrelations. One such interrelation is the existence of various dualities connecting differently looking models. Another one is a rather mysterious appearance of classical integrable equations as exact relations built into the structure of quantum integrable systems even at $\hbar\neq 0$. Moreover, such relations sometimes allow one to develop alternative methods to diagonalize commuting Hamiltonians of the quantum systems, based on purely classical theory.
The latter program was first realized in [2] (see also [3, 4]), where functional relations for higher conserved quantities of generalized quantum spin chains [5, 6, 7] were interpreted as a classical discrete dynamical system which was identified with the Hirota bilinear difference equation [8] known in soliton theory since 1981. (It is often referred to as a fully discrete KP equation.) It was demonstrated that this interpretation allows one to solve the spectral problem for spin chains using methods of classical soliton theory. Later, in [9, 10] (see also [11] for a further development) this approach was extended to generalized spin chains with superalgebra symmetries (super-spin chains or “graded magnets” introduced by in [12]). This extension was based on the functional relations for higher transfer matrices of the super-spin chains established in [13, 14].
A deeper understanding of the connection between quantum spin chains and soliton equations was achieved in the works [15][22], where continuous flows parametrized by infinitely many continuous parameters ${\bf t}=\{t_{1},t_{2},t_{3},\ldots\}$ (“time variables”) were defined in the space of commuting conserved quantities of a quantum spin chain. The dynamics in the times ${\bf t}$ was then identified with the modified Kadomtsev-Petviashvili (mKP) hierarchy of nonlinear integrable equations. In this paper we give a review of the works [15][22], refining some arguments from them and making some statements more precise.
[…]
${\sf R}^{\lambda}_{01}(x)=x{\sf I}+\eta\sum_{a,b=1}^{n}\pi_{\lambda}({\bf e}_{ab})^{(0)}\otimes e_{ba}^{(1)}$ (1.1)
acting in the tensor product of two linear spaces $V_{0}\times V_{1}$, where ${\sf I}$ is the unity matrix, ${\bf e}_{ab}$ are generators of the universal enveloping algebra $U(gl_{n})$, $\pi_{\lambda}$ is the finite-dimensional irreducible representation of $U(gl_{n})$ associated with the Young diagram $\lambda$ and $e_{ab}$ is ${\bf e}_{ab}$ in the $n$-dimensional vector representation. The transfer matrix ${\sf T}_{\lambda}(x)$ is defined as
${\sf T}_{\lambda}(x)=\mbox{tr}_{0}\Bigl{(}{\sf R}^{\lambda}_{01}(x-x_{1}){\sf R}^{\lambda}_{02}(x-x_{2})\ldots{\sf R}^{\lambda}_{0N}(x-x_{N})\,\pi_{\lambda}({\bf g}_{0})\Bigr{)},$ (1.2)
where ${\bf g}=\mbox{diag}(p_{1},\ldots,p_{n})$ is the diagonal twist matrix and $x_{i}$ are arbitrary (in general, complex) numbers called inhomogeneity parameters (we assume that all of them are distinct). The standard method of simultaneous diagonalization of the transfer matrices is the Bethe ansatz, coordinate or algebraic [23][26]. For $n>2$ this technique is called the nested Bethe ansatz. As a result, the eigenvalues of the transfer matrices are expressed through a set of auxiliary quantities (Bethe roots) satisfying a system of algebraic equations called (nested) Bethe equations (see [27, 28]).
Let ${\sf T}(x,{\bf t})$ be the following generating function of the commuting transfer matrices:
${\sf T}(x,{\bf t})=\sum_{\lambda}{\sf T}_{\lambda}(x)s_{\lambda}({\bf t}),$ (1.3)
where $s_{\lambda}({\bf t})$ are Schur polynomials [29] and the sum goes over all Young diagrams including the empty one (actually, for $GL(n)$-invariant models the sum is restricted to diagrams with not more than $n$ non-empty rows). In [15] the generating function (1.3) was called the master $T$-operator (see also [30], where it was introduced in an implicit form). Its main property, proved in [15], is that it satisfies the bilinear equations for the tau-function of the mKP hierarchy, with $x$ playing the role of the “zeroth time” variable. The generating bilinear equation has the form [31, 32]
[…]
A further analysis shows that the objects arising in the process of solving the quantum model using the algebraic Bethe ansatz technique and specific for the quantum theory have their counterparts in the classical theory of soliton equations. For example, the classical facet of the nested Bethe ansatz procedure is a chain of Bäcklund transformations (which we call the “undressing chain”) that gradually “undress” a solution of the mKP hierarchy to the trivial one. Each step of this chain corresponds to a level of the nested Bethe ansatz. The eigenvalues of the Baxter’s $Q$-operators are nothing else than tau-functions arising at different steps of the undressing chain. Then the Bethe equations for their roots acquire a nice interpretation of equations of motion for the Ruijsenaars-Schneider system of particles in discrete time (each step of the discrete time corresponds to a level of the Bethe ansatz). Further, the non-commutative generating function for transfer matrices in fundamental representations becomes, on the classical side, the so-called wave operator ${\bf W}(x)$ (known also as the dressing operator). In general, it is an infinite series in inverse powers of the shift operator $e^{\eta\partial_{x}}$ but a characteristic feature of solutions relevant to the $GL(n)$-invariant quantum spin chains is that this series truncates at the $n$-th term, and this operator admits a factorization as a finite product of first order difference operators. This factorization plays a key role in the algebraic Bethe ansatz solution since the coefficients of the first order operators in each factor are constructed from eigenvalues of the $Q$-operators $Q_{k}(x)$, $k=1,\ldots,n-1$, so eigenvalues of the transfer matrices can be expressed through the $Q_{k}$’s whose roots are subject to the system of Bethe equations.
The mKP hierarchy has a lot of solutions of very different nature. So, to make the correspondence with the classical theory complete, one should characterize the class of solutions corresponding to eigenvalues of the master $T$-operator. Since its matrix elements are polynomial in $x$ of degree $N$ and the master $T$-operators for all $x$ can be simultaneously diagonalized, the same is true for the eigenvalues, i.e., the eigenvalues $T(x,{\bf t})$ are polynomials in $x$ of degree $N$ whose roots depend on the times $t_{i}$:
$T(x;{\bf t})=C({\bf t})\prod_{k=1}^{N}\left(x-x_{k}({\bf t})\right).$ (1.5)
The dynamics of the zeros of polynomial tau-functions is a well known subject in the theory of integrable nonlinear partial differential equations. In the works of Krichever and others (see [33][40]) it was found that these dynamics are described by equations of motion of integrable many-body systems of particles of the Calogero-Moser and Ruijsenaars-Schneider type. In particular, the dynamics of zeros of the tau-function of the mKP hierarchy of the form (1.5) in the times $t_{k}$ coincide with the dynamics of the Ruijsenaars-Schneider system of particles [41] (which is also known as a relativistic deformation of the Calogero-Moser system [42, 43], the parameter $\eta$ playing the role of the inverse velocity of light) with respect to the $k$-th Hamiltonian flow (for all $k$ this result was obtained in [44]). The Ruijsenaars-Schneider system admits the Lax representation with the Lax matrix
$L_{ij}=\frac{\partial_{t_{1}}x_{i}}{x_{i}-x_{j}-\eta}.$ (1.6)
The time evolution is an isospectral transformation of $L$ and the characteristic polynomial $\det(zI-L)$ is a generating function of conserved quantities.
From this it follows a nontrivial connection between the generalized inhomogeneous quantum spin chains solvable by the algebraic Bethe ansatz and classical integrable many-body systems of the Calogero-Moser type. This connection is called the quantum-classical duality. To formulate it, we need to consider the transfer matrix ${\sf T}_{(1)}(x)$ (1.2) corresponding to the vector representation (for which $\lambda$ is one box) and define quantum commuting Hamiltonians of the inhomogeneous spin chain as
${\bf H}_{i}=\eta^{-1}\mathop{\hbox{res}}\limits_{x=x_{i}}\left(\frac{{\sf T}_{(1)}(x)}{\prod\limits_{k=1}^{N}(x-x_{k})}\right).$ (1.7)
Then the quantum-classical duality states that the eigenvalues $H_{i}$ of the ${\bf H}_{i}$’s are given by velocities $\dot{x}_{i}(0)=\partial_{t_{1}}x_{i}\Bigr{|}_{{\bf t}=0}$ of the classical Ruijsenaars-Schneider particles at ${\bf t}=0$:
$H_{i}=-\eta^{-1}\dot{x}_{i}(0),$ (1.8)
which have to be found from the condition that the spectrum of the Lax matrix $L$ has the form
$\mbox{Spec}\,L=\Bigl{(}\underbrace{p_{1},\ldots,p_{1}}_{M_{1}},\,\underbrace{p_{2},\ldots,p_{2}}_{M_{2}},\,\ldots,\,\underbrace{p_{n},\ldots,p_{n}}_{M_{n}}\Bigr{)},$ (1.9)
where $M_{a}$ are the eigenvalues of the weight operators ${{\bf M}_{a}=\sum_{k=1}^{N}\pi_{(1)}({\bf e}^{(k)}_{aa})}$ on the common eigenstate with the transfer matrix and $p_{i}$ are the twist parameters (elements of the diagonal twist matrix ${\bf g}$).
A direct proof of this statement was given in [16] for models with rational dependence on the spectral parameter $x$ and in [19] for models with trigonometric $R$-matrices. By other reasoning and in another context, this kind of duality was discussed in [45] (see also [46, 47], where similar phenomena were observed in some simpler particular and limiting cases). However, the proof given in [16] was essentially based on the nested Bethe equations and was technically involved. In this paper, we suggest another, much simpler proof, which avoids any explicit use of Bethe equations.
The structure of the paper is as follows. In Section 2 we recall the standard definitions and facts from the theory of integrable $GL(n)$-invariant spin chains. In Section 3 we introduce the higher transfer matrices ${\sf T}_{\lambda}(x)$ indexed by the Young diagrams $\lambda$ and recall the functional relations for them, which have the form of the spectral parameter dependent determinant identities of the Jacobi-Trudi type. The master $T$-operator is introduced in Section 4.1, and in Section 4.2 the integral bilinear relation for it is given. This relation allows one to identify eigenvalues of the master $T$-operator with tau-functions of the mKP hierarchy. Section 5 is devoted to the basic facts related to the mKP hierarchy. In particular, we recall the construction of the wave operator ${\bf W}(x,{\bf t})$, the wave function $\psi(x,{\bf t};z)$ and its adjoint $\psi^{*}(x,{\bf t};z)$. In Section 6 we study (quasi)polynomial solutions to the mKP hierarchy. They are constructed using the approach suggested by Krichever, which is based on the conditions (6.2) imposed on the wave function as a function of the spectral parameter $z$. Also, in this section the undressing and dressing Bäcklund transformations are considered, and a chain of the Bäcklund transformations is introduced. The index that labels steps of the chain can be regarded as a discrete time variable. The dynamics of zeros of the tau-functions in this discrete time is the Ruijsenaars-Schneider system in discrete time. Remarkably, its equations of motion are the nested Bethe equations. In Section 7 we present a detailed identification of the objects from the algebraic Bethe ansatz with the ones from the classical theory of the mKP hierarchy. In Section 8 the connection with the classical Ruijsenaars-Schneider many-body system is established, and the quantum-classical duality is discussed. Finally, in Section 9 we make a few concluding remarks and mention some open problems.
2 Generalized spin chains
2.1 $GL(n)$-invariant $R$-matrices
[…]
Let $V_{1},V_{2},V_{3}$ be three copies of the space $V$. By ${\sf R}_{12}(x)$ we denote the $R$-matrix that acts nontrivially on $V_{1}\otimes V_{2}$ and trivially in $V_{3}$, etc. The $R$-matrix is required to satisfy the Yang-Baxter equation
${\sf R}_{12}(x-x^{\prime}){\sf R}_{13}(x){\sf R}_{23}(x^{\prime})={\sf R}_{23}(x^{\prime}){\sf R}_{13}(x){\sf R}_{12}(x-x^{\prime}),$ (2.1)
where the both sides are linear operators on the space $V_{1}\otimes V_{2}\otimes V_{3}$. The $GL(n)$-invariance means that the ${\sf R}(x)$ commutes with ${\bf g}\otimes{\bf g}$ for any ${\bf g}\in GL(n)$, i.e.,
${\bf g}\otimes{\bf g}\,{\sf R}(x)={\sf R}(x)\,{\bf g}\otimes{\bf g}\qquad\mbox{or}\qquad{\bf g}_{1}{\bf g}_{2}{\sf R}_{12}(x)={\sf R}_{12}(x){\bf g}_{1}{\bf g}_{2}.$ (2.2)
It is known that there are $GL(n)$-invariant solutions of the Yang-Baxter equation of size $n^{2}\!\times\!n^{2}$ for any $n\geq 2$ having polynomial dependence on the spectral parameter. These $R$-matrices have the form
${\sf R}(x)=x{\sf I}+\eta\sum_{a,b=1}^{n}e_{ab}\otimes e_{ba},$ (2.3)
where ${\sf I}$ is the unity matrix in $V\otimes V$, $e_{ab}$ are elementary $n\!\times\!n$ matrices with 1 at the place $ab$ and 0 otherwise and $\eta$ is a parameter [^1] This parameter can be absorbed in $x$ but we prefer to keep it alive in order to prepare for the limit to the Gaudin model which is $\eta\to 0$.. Note that
${\sf P}=\sum_{a,b=1}^{n}e_{ab}\otimes e_{ba}$
is the permutation operator in $V\otimes V$, so the $R$-matrix can be written as ${\sf R}(x)=x{\sf I}+\eta{\sf P}$. In what follows, the permutation operator in $V_{i}\otimes V_{j}$ is denoted by ${\sf P}_{ij}$.
2.2 Transfer matrices
[…]
${\sf T}(x)=\mbox{tr}_{0}\Bigl{(}{\sf R}_{01}(x-x_{1}){\sf R}_{02}(x-x_{2})\ldots{\sf R}_{0N}(x-x_{N})\Bigr{)},$
where the trace is taken in the auxiliary space $V_{0}$. The transfer matrix is a linear operator in the space ${\cal V}$. The Yang-Baxter equation for the $R$-matrix guarantees that the transfer matrices commute at different values of the spectral parameter: $[{\sf T}(x),\,{\sf T}(x^{\prime})]=0.$
More generally, one can also consider the chain with quasiperiodic (twisted) boundary conditions inserting under the trace a group element ${\bf g}\in GL(n)$ (twist), which for simplicity we assume to be diagonal (${\bf g}=\mbox{diag}(g_{1},g_{2},\ldots,g_{n})$):
${\sf T}(x)=\mbox{tr}_{0}\Bigl{(}{\sf R}_{01}(x-x_{1}){\sf R}_{02}(x-x_{2})\ldots{\sf R}_{0N}(x-x_{N})\,{\bf g}_{0}\Bigr{)}$ (2.4)
The notation ${\bf g}_{0}$ means that ${\bf g}$ acts in the auxiliary space (number 0). The $GL(n)$-invariance (2.2) implies that these transfer matrices commute at different values of the spectral parameter. In the homogeneous chain with periodic boundary conditions (at $x_{j}=0$, ${\bf g}={\sf I}$) there exists a local Hamiltonian which is the logarithmic derivative of ${\sf T}(x)$ at $x=0$. In inhomogeneous chains local Hamiltonians commuting with the transfer matrix in general do not exist.
Matrix elements of the transfer matrix (2.4) are polynomials in $x$ of degree not greater than $N$. Let us normalize the transfer matrix in a different way, dividing it by the polynomial
$\phi(x)=\prod_{j=1}^{N}(x-x_{j}).$ (2.5)
In this normalization the transfer matrix
${\bf T}(x)=\frac{{\sf T}(x)}{\prod\limits_{j=1}^{N}(x-x_{j})}$ (2.6)
has simple poles at the points $x_{j}$. Obviously, the transfer matrix ${\bf T}(x)$ is given by
${\bf T}(x)=\mbox{tr}_{0}\Bigl{(}\tilde{\sf R}_{01}(x-x_{1})\tilde{\sf R}_{02}(x-x_{2})\ldots\tilde{\sf R}_{0N}(x-x_{N})\,{\bf g}_{0}\Bigr{)},$ (2.7)
where
${\tilde{\sf R}(x)={\sf I}+\frac{\eta}{x}\,{\sf P}}$
is the $R$-matrix which differs from the ${\sf R}(x)$ by a scalar factor. One can introduce Hamiltonians ${\bf H}_{j}$ of the inhomogeneous spin chain as residues of ${\bf T}(x)$ at the poles:
${\bf T}(x)=\mbox{tr}\,{\bf g}+\sum_{j=1}^{N}\frac{\eta{\bf H}_{j}}{x-x_{j}}.$ (2.8)
These operators commute with each other. However, they are non-local. Their explicit form is as follows:
${\bf H}_{i}=\tilde{\sf R}_{i\,i\!-\!1}(x_{i}-x_{i-1})\ldots\tilde{\sf R}_{i1}(x_{i}-x_{1}){\bf g}_{i}\tilde{\sf R}_{iN}(x_{i}-x_{N})\ldots\tilde{\sf R}_{i\,i\!+\!1}(x_{i}-x_{i+1}).$ (2.9)
Comparing the expansions as $x\to\infty$ of (2.8) and
$\begin{array}[]{lll}{\bf T}(x)&=&\displaystyle{\mbox{tr}_{0}\left[\Bigl{(}{\sf I}+\frac{\eta{\sf P}_{01}}{x-x_{1}}\Bigr{)}\ldots\Bigl{(}{\sf I}+\frac{\eta{\sf P}_{0N}}{x-x_{N}}\Bigr{)}\,{\bf g}_{0}\right]}\\ &&\\ &=&\displaystyle{\mbox{tr}\,{\bf g}\cdot{\sf I}+\frac{\eta}{x}\sum_{i=1}^{N}\mbox{tr}_{0}\Bigl{(}{\sf P}_{0i}{\bf g}_{0}\Bigr{)}+\ldots=\mbox{tr}\,{\bf g}\cdot{\sf I}+\frac{\eta}{x}\sum_{i=1}^{N}{\bf g}_{i}+\ldots,}\end{array}$
we get the following “sum rule”:
$\sum_{i=1}^{N}{\bf H}_{i}=\sum_{j=1}^{N}{\bf g}_{j},$
where both sides are operators on ${\cal V}=V_{1}\otimes V_{2}\otimes\ldots\otimes V_{N}$ and ${\bf g}_{j}$ acts as ${\bf g}$ in $V_{j}$ and trivially in the other tensor factors.
Let us mention the limit of this construction as $\eta\to 0$, which is of its own interest. In this limit the generalized spin chain becomes the Gaudin model. Set ${\bf g}=e^{\eta{\bf h}}$, then in the limit $\eta\to 0$ we have ${\bf H}_{i}=1+\eta{\bf H}^{G}_{i}+O(\eta^{2})$, where
${\bf H}^{G}_{i}={\bf h}_{i}+\sum_{j\neq i}\frac{{\sf P}_{ij}}{x_{i}-x_{j}}$
are commuting Gaudin Hamiltonians of the Gaudin model.
Let $e_{ab}^{(j)}$ be the operator in ${\cal V}$ that acts as $e_{ab}$ in $V_{j}$ and trivially in the other tensor factors. Consider the operators
${\bf M}_{a}=\sum_{j=1}^{N}e_{aa}^{(j)},\qquad a=1,\ldots,n$ (2.10)
which are sometimes called weight operators [^2] In the case $n=2$ which corresponds to the XXX spin chain with spins $\frac{1}{2}$ the operator $\frac{1}{2}({\bf M}_{1}-{\bf M}_{2})$ is the operator of the $z$-projection of the total spin.. It is easy to see that they commute with the transfer matrix and among themselves and can be simultaneously diagonalized. Therefore, one can find eigenstates of the transfer matrix which are simultaneously eigenstates of the operators ${\bf M}_{a}$ with eigenvalues $M_{a}$. Note that ${\sum_{a=1}^{n}{\bf M}_{a}=N\,{\sf I}}$, so ${\sum_{a=1}^{n}M_{a}=N}$.
Let us present the result of diagonalization of the transfer matrix ${\bf T}(x)$ by means of the algebraic (nested) Bethe ansatz. We give it here without derivation (see [27, 28] for details). The eigenvalues $\Lambda(x)$ of ${\bf T}(x)$ are given by
$\Lambda(x)=\sum_{b=1}^{n}g_{b}\prod_{\gamma=1}^{{\cal N}_{b-1}}\frac{x-w_{\gamma}^{(b-1)}+\eta}{x-w_{\gamma}^{(b-1)}}\prod_{\beta=1}^{{\cal N}_{b}}\frac{x-w_{\beta}^{(b)}-\eta}{x-w_{\beta}^{(b)}},$ (2.11)
where ${\cal N}_{0}=N$, ${\cal N}_{0}\geq{\cal N}_{1}\geq{\cal N}_{2}\geq\ldots\geq{\cal N}_{n-1}\geq 0$ are non-negative integers, ${\cal N}_{n}=0$, $w_{\gamma}^{(0)}=x_{\gamma}$ and the sets of Bethe roots $\{w_{\beta}^{(b)}\}_{\beta=1}^{{\cal N}_{b}}$ satisfy the system of nested Bethe ansatz equations
$g_{b}\prod_{\gamma=1}^{{\cal N}_{b-1}}\frac{w_{\alpha}^{(b)}-w_{\gamma}^{(b-1)}+\eta}{w_{\alpha}^{(b)}-w_{\gamma}^{(b-1)}}=g_{b+1}\prod_{\gamma\neq\alpha}^{{\cal N}_{b}}\frac{w_{\alpha}^{(b)}-w_{\gamma}^{(b)}+\eta}{w_{\alpha}^{(b)}-w_{\gamma}^{(b)}-\eta}\prod_{\beta=1}^{{\cal N}_{b+1}}\frac{w_{\alpha}^{(b)}-w_{\beta}^{(b+1)}-\eta}{w_{\alpha}^{(b)}-w_{\beta}^{(b+1)}}.$ (2.12)
Here $b=1,\ldots,n-1$, $\alpha=1,\ldots,{\cal N}_{b}$. The numbers ${\cal N}_{a}$ are such that ${\cal N}_{a-1}-{\cal N}_{a}=M_{n-a+1}$, $a=1,\ldots,n$, where $M_{a}$ are eigenvalues of the operators ${\bf M}_{a}$. The total number of equations in the system is ${\cal N}_{1}+\ldots+{\cal N}_{n-1}$. As it follows from (2.8), (2.11), the eigenvalues $H_{i}$ of the Hamiltonians ${\bf H}_{i}$ are given by
$H_{i}=g_{1}\prod_{k\neq i}^{N}\frac{x_{i}-x_{k}+\eta}{x_{i}-x_{k}}\prod_{\gamma=1}^{{\cal N}_{1}}\frac{x_{i}-w_{\gamma}^{(1)}-\eta}{x_{i}-w_{\gamma}^{(1)}}.$ (2.13)
In Section 6.4 we will obtain the system of Bethe equations in the context of the mKP hierarchy.
3 Transfer matrices as generalized characters
For $GL(n)$-invariant models with $n>2$ the algebra of commuting operators is larger than the one generated by the Hamiltonians ${\bf H}_{j}$. In fact, there exists a larger family of (more general) transfer matrices commuting with ${\sf T}(x)$.
To proceed, we need some information about representations of the group $GL(n)$ and the universal enveloping algebra $U(gl_{n})$ which has generators ${\bf e}_{ab}$ with the commutation relations
${\bf e}_{ab}{\bf e}_{a^{\prime}b^{\prime}}-{\bf e}_{a^{\prime}b^{\prime}}{\bf e}_{ab}=\delta_{a^{\prime}b}{\bf e}_{ab^{\prime}}-\delta_{ab^{\prime}}{\bf e}_{a^{\prime}b}.$ (3.1)
[…]
Two simple special cases are important for what follows. The empty diagram $\lambda=\emptyset$ corresponds to the trivial representation: $\pi_{\emptyset}({\bf e}_{ab})=0$. The column of $n$ boxes corresponds to the one-dimensional representation $\pi_{(1^{n})}({\bf e}_{ab})=\delta_{ab}$.
3.1 Higher transfer matrices
[…]
${\sf R}^{\lambda}(x)=x{\sf I}+\eta\sum_{a,b}\pi_{\lambda}({\bf e}_{ab})\otimes e_{ba}.$ (3.2)
The $GL(n)$-invariance means that
$\pi_{\lambda}({\bf g})\otimes{\bf g}\,{\sf R}^{\lambda}(x)={\sf R}^{\lambda}(x)\,\pi_{\lambda}({\bf g})\otimes{\bf g}.$
The $R$-matrices ${\sf R}^{\lambda}(x)$ satisfy the Yang-Baxter equation
${\sf R}^{\lambda\mu}_{12}(x-x^{\prime}){\sf R}^{\lambda}_{13}(x){\sf R}^{\mu}_{23}(x^{\prime})={\sf R}^{\mu}_{23}(x^{\prime}){\sf R}^{\lambda}_{13}(x){\sf R}^{\lambda\mu}_{12}(x-x^{\prime}),$ (3.3)
where ${\sf R}^{\lambda\mu}(x-x^{\prime})$ is some $R$-matrix acting in the tensor product $V_{\lambda}\otimes V_{\mu}$. Its explicit form for arbitrary $\lambda,\mu$ is complicated.
[…]
${\sf T}_{\lambda}(x)=\mbox{tr}_{V_{\lambda}}\Bigl{(}{\sf R}^{\lambda}_{01}(x-x_{1}){\sf R}^{\lambda}_{02}(x-x_{2})\ldots{\sf R}^{\lambda}_{0N}(x-x_{N})\,\pi_{\lambda}({\bf g}_{0})\Bigr{)}.$ (3.4)
From the Yang-Baxter equation (3.3) and $GL(n)$-invariance it follows that the transfer matrices ${\sf T}_{\lambda}(x)$ commute for different $x$ and $\lambda$:
$[{\sf T}_{\lambda}(x),\,{\sf T}_{\mu}(x^{\prime})]=0.$
As it was already mentioned, the empty diagram $\lambda=\emptyset$ corresponds to the trivial representation $\pi_{\emptyset}({\bf e}_{ab})=0$, $\pi_{\emptyset}({\bf g})=1$ and we have from (3.4):
${\sf T}_{\emptyset}(x)=\prod_{i=1}^{N}(x-x_{i})\cdot{\sf I}=\phi(x){\sf I}.$ (3.5)
One can introduce normalized transfer matrices ${\bf T}_{\lambda}(x)$ dividing by ${\sf T}_{\emptyset}(x)$:
${\bf T}_{\lambda}(x)=\frac{{\sf T}_{\lambda}(x)}{{\sf T}_{\emptyset}(x)}.$
Then ${\bf T}_{\emptyset}(x)={\sf I}$ and ${\bf T}_{(1)}(x)={\bf T}(x)$ introduced in (2.7).
For the one-column diagram $\lambda=(1^{n})$ of height $n$ we have one-dimensional representation $\pi_{(1^{n})}({\bf e}_{ab})=\delta_{ab}$, $\pi_{(1^{n})}({\bf g})=\det{\bf g}$ and formula (3.4) yields:
${\sf T}_{(1^{n})}(x)=\det{\bf g}\,\phi(x+\eta)\,{\sf I}.$ (3.6)
In the quantum inverse scattering method this transfer matrix has the meaning of quantum determinant of the quantum monodromy matrix. For diagrams such that $\ell(\lambda)>n$ the transfer matrices vanish identically.
At $N=0$ the definition (3.4) yields:
${\bf T}^{(N=0)}_{\lambda}(x)=\mbox{tr}_{V_{\lambda}}\pi_{\lambda}({\bf g})=\chi_{\lambda}({\bf g}),$ (3.7)
where $\chi_{\lambda}({\bf g})$ is the character of ${\bf g}$ in the representation $\pi_{\lambda}$. Also, we have
${\bf T}_{\lambda}(x)=\chi_{\lambda}({\bf g})\cdot{\sf I}+O(1/x),\qquad x\to\infty,$
so the normalized transfer matrices can be regarded as a generalization of characters.
It is known that the characters are given by Schur polynomials $s_{\lambda}$ of the eigenvalues $g_{i}$ of the matrix ${\bf g}$:
$\chi_{\lambda}({\bf g})=s_{\lambda}(\{g_{i}\})=\frac{\det_{ij}\left(g_{i}^{n+\lambda_{j}-j}\right)}{\det_{ij}\left(g_{i}^{n-j}\right)}.$
The Schur polynomials are symmetric functions of $g_{i}$. It is often convenient to consider Schur polynomials $s_{\lambda}(\{\xi_{i}\})$, where $\{\xi_{i}\}$ is a set of some variables, as functions $s_{\lambda}({\bf t})$ of their power sums $t_{k}=\frac{1}{k}\sum\limits_{i}\xi_{i}^{k}$ (${\bf t}=\{t_{1},t_{2},t_{3},\ldots\}$ is the set of these new variables, in general infinite). For example,
$\begin{array}[]{c}s_{\emptyset}({\bf t})=1,\quad s_{(1)}({\bf t})=t_{1},\quad s_{(2)}({\bf t})=\frac{1}{2}t_{1}^{2}+t_{2},\quad s_{(1^{2})}({\bf t})=\frac{1}{2}t_{1}^{2}-t_{2}\end{array}$
and so on. For any finite diagram $\lambda$ the polynomial $s_{\lambda}({\bf t})$ depends only on a finite number of $t_{i}$’s.
The Schur polynomials corresponding to the diagrams that are rows or columns of the form, respectively, $\lambda=(s)$ or $\lambda=1^{a}$ play a specially important role. It is customary to use the special notation for them [29]:
$h_{k}({\bf t})=s_{(k)}({\bf t}),\qquad e_{k}({\bf t})=s_{(1^{k})}({\bf t}).$ (3.8)
The generating functions for them are as follows:
$\begin{array}[]{l}\displaystyle{\exp\Bigl{(}\sum_{k\geq 1}t_{k}z^{k}\Bigr{)}=\sum_{k\geq 0}h_{k}({\bf t})z^{k},}\\ \\ \displaystyle{\exp\Bigl{(}-\sum_{k\geq 1}t_{k}(-z)^{k}\Bigr{)}=\sum_{k\geq 0}e_{k}({\bf t})z^{k}.}\end{array}$ (3.9)
After the substitution $t_{k}=\frac{1}{k}\sum\limits_{i=1}^{m}\xi_{i}^{k}$ these polynomials become symmetric functions of the variables $\xi_{1},\ldots,\xi_{m}$. Note that $e_{m}(\{\xi_{i}\}_{m})=\xi_{1}\ldots\xi_{m}$ and $e_{k}(\{\xi_{i}\}_{m})=0$ if $m>k$. From (3.9) it is clear that
$e_{k}({\bf t})=(-1)^{k}h_{k}(-{\bf t}).$ (3.10)
The Schur polynomials satisfy a number of important identities. First, we should mention the Cauchy-Littlewood identity [29]
$\sum_{\lambda}s_{\lambda}({\bf t})s_{\lambda}({\bf t}^{\prime})=\exp\Bigl{(}\sum_{k\geq 1}kt_{k}t^{\prime}_{k}\Bigr{)},$ (3.11)
where the sum on the left-hand side is taken over all Young diagrams including the empty one. Second, there are identities which express Schur polynomials for general $\lambda$’s through those corresponding to the diagrams which are either rows or columns, i.e., through $h_{\lambda}({\bf t})$ or $e_{\lambda}({\bf t})$:
$s_{\lambda}({\bf t})=\det_{1\leq i,j\leq\lambda_{1}^{\prime}}h_{\lambda_{i}-i+j}({\bf t}),$ (3.12)
$s_{\lambda}({\bf t})=\det_{1\leq i,j\leq\lambda_{1}}e_{\lambda_{i}^{\prime}-i+j}({\bf t}).$ (3.13)
Here $\lambda^{\prime}$ is the diagram $\lambda$ transposed with respect to the main diagonal, so that $\lambda_{1}^{\prime},\lambda_{2}^{\prime},\ldots$ are heights of columns of $\lambda$. Equations (3.12), (3.13) are called the Jacobi-Trudi identities. In terms of characters, they have the form
$\chi_{\lambda}({\bf g})=\det_{1\leq i,j\leq\lambda_{1}^{\prime}}\chi_{(\lambda_{i}-i+j)}({\bf g}),$ (3.14)
$\chi_{\lambda}({\bf g})=\det_{1\leq i,j\leq\lambda_{1}}\chi_{(1^{\lambda_{i}^{\prime}-i+j})}({\bf g}).$ (3.15)
Note that $\chi_{\lambda}({\bf g})=0$ if $\ell(\lambda)>n$.
The transfer matrices in symmetric or antisymmetric representations, i.e., corresponding to the diagrams which are rows or columns of the form, respectively, $\lambda=(s)$ or $\lambda=1^{a}$ play a specially important role. In what follows we will use the special simplified notation for them:
${\sf T}_{s}(x)={\sf T}_{(s)}(x),\qquad{\sf T}^{a}(x)={\sf T}_{(1^{a})}(x).$ (3.16)
The analogy between transfer matrices and characters is further supported by the fact that the transfer matrices satisfy the following identities (functional relations), which look similarly to the Jacobi-Trudi identities for characters:
${\bf T}_{\lambda}(x)=\det_{1\leq i,j\leq\lambda_{1}^{\prime}}{\bf T}_{\lambda_{i}-i+j}(x-(j\!-\!1)\eta),$ (3.17)
${\bf T}_{\lambda}(x)=\det_{1\leq i,j\leq\lambda_{1}}{\bf T}^{\lambda_{i}^{\prime}-i+j}(x+(j\!-\!1)\eta).$ (3.18)
They are called the Cherednik-Bazhanov-Reshetikhin (CBR) identities or quantum Jacobi-Trudi identities [5, 6]. For the transfer matrices in the initial normalization (3.4), they look as follows:
${\sf T}_{\lambda}(x)=\left(\prod_{k=1}^{\lambda_{1}^{\prime}-1}\phi(x-k\eta)\right)^{-1}\det_{1\leq i,j\leq\lambda_{1}^{\prime}}{\sf T}_{\lambda_{i}-i+j}(x-(j\!-\!1)\eta),$ (3.19)
${\sf T}_{\lambda}(x)=\left(\prod_{k=1}^{\lambda_{1}-1}\phi(x+k\eta)\right)^{-1}\det_{1\leq i,j\leq\lambda_{1}}{\sf T}^{\lambda_{i}^{\prime}-i+j}(x+(j\!-\!1)\eta).$ (3.20)
In fact the relations (3.19) and (3.20) are equivalent: (3.20) follows from (3.19) and vice versa. Note that since ${\sf T}_{\lambda}(x)$ are polynomials, the possible poles at $x=x_{i}\pm k\eta$ coming from the pre-factors in the right-hand sides, must be cancelled by zeros of the determinants. Like the characters, the transfer matrices ${\sf T}_{\lambda}(x)$ vanish identically if $\lambda_{1}^{\prime}>n$ (in particular, ${\sf T}^{a}(x)$ vanishes if $a>n$).
3.2 The approach based on co-derivative
There is an elegant way, suggested in [48], to represent the transfer matrices ${\sf T}_{\lambda}(x)$ through matrix derivatives of the characters $\chi_{\lambda}({\bf g})$ with respect to the twist matrix ${\bf g}$ (for this purpose it is not assumed to be diagonal). This approach is an alternative to the fusion procedure.
Let $f({\bf g})$ be any function on the group $GL(n)$ (${\bf g}\in GL(n)$). Define the matrix derivative (which is called co-derivative in [48]) as follows:
$Df({\bf g})=\sum_{a,b}e_{ab}\frac{\partial}{\partial\varepsilon}f\left(e^{\varepsilon{\bf e}_{ba}}{\bf g}\right)\Bigr{|}_{\varepsilon=0}.$ (3.21)
[…]
$D\det{\bf g}=\det{\bf g}\cdot{\sf I}.$ (3.22)
An equivalent definition in components is
$D_{b}^{a}=\sum_{c}g^{a}_{c}\frac{\partial}{\partial g^{b}_{c}},$
where $g^{a}_{b}$ are matrix elements of the matrix ${\bf g}\in GL(n)$ in the vector representation. Explicitly, we have:
$D_{b}^{a}f({\bf g})=\frac{\partial}{\partial\varepsilon}f\left(e^{\varepsilon{\bf e}_{ba}}{\bf g}\right)\Bigr{|}_{\varepsilon=0}.$
A direct calculation of the commutator $[D_{b_{2}}^{a_{2}},D_{b_{1}}^{a_{1}}]$ shows that
$[D_{b_{2}}^{a_{2}},D_{b_{1}}^{a_{1}}]=\delta_{a_{1}b_{2}}D^{a_{2}}_{b_{1}}-\delta_{a_{2}b_{1}}D^{a_{1}}_{b_{2}},$ (3.23)
i.e., the operators $D^{a}_{b}$ have the same commutation relations as the generators ${\bf e}_{ab}$ of the algebra $U(gl_{n})$.
In the case when the co-derivatives act on functions with values in $\mbox{End}(\otimes_{i}V_{i})$, it is convenient to modify the notation by adding index $i$ which numbers the spaces in the tensor product:
$D_{i}f({\bf g})=\sum_{a,b}e_{ab}^{(i)}\frac{\partial}{\partial\varepsilon}f\left(e^{\varepsilon{\bf e}_{ba}}{\bf g}\right)\Bigr{|}_{\varepsilon=0}.$
Here $e_{ab}^{(i)}$ acts non-trivially in $V_{i}$. In this notation we have, for example: $D_{1}\mbox{tr}\,{\bf g}={\bf g}_{1}$, $D_{2}{\bf g}_{1}={\sf P}_{21}{\bf g}_{1}$, while the relation (3.23) is written in the form $[D_{2},D_{1}]={\sf P}_{12}(D_{1}-D_{2})$.
A careful analysis shows that the transfer matrix ${\sf T}_{\lambda}(u)$ can be expressed as
${\sf T}_{\lambda}(x)=(x-x_{N}+\eta D_{N})\ldots(x-x_{1}+\eta D_{1})\chi_{\lambda}({\bf g}).$ (3.24)
With the help of this representation, the authors of [48] managed to prove the CBR identities for models with rational $R$-matrices in a direct way. Here we only mention that this representation allows one to obtain the expression for the quantum determinant in a very easy way: in this case $\chi_{(1^{n})}({\bf g})=\det{\bf g}$, and we obtain
${\sf T}^{n}(x)=\det{\bf g}\prod_{j=1}^{N}(x-x_{j}+\eta)\cdot{\sf I},$ (3.25)
where we have used (3.22).
4 The master $T$-operator as an operator-valued tau-function
4.1 The master $T$-operator
It turns out to be instructive to consider a generating function for the transfer matrices ${\sf T}_{\lambda}(x)$. As such, it was introduced in [15] under the name of master $T$-operator (see also [30] for its preliminary version). Let ${\bf t}=\{t_{1},t_{2},t_{3},\ldots\}$ be an infinite set of complex variables. The master $T$-operator is defined as an infinite series of the form
${\sf T}(x;{\bf t})=\sum_{\lambda}s_{\lambda}({\bf t}){\sf T}_{\lambda}(x),$ (4.1)
where the sum, like in (3.11), is taken over all Young diagrams including the empty one. Since ${\sf T}_{\lambda}(x)=0$ if $\lambda_{1}^{\prime}>n$, the sum in (4.1) goes over the diagrams with $\lambda_{1}^{\prime}\leq n$. Like ${\sf T}_{\lambda}(x)$, ${\sf T}(x;{\bf t})$ is an operator in the space ${\cal V}$. It depends on the elements $g_{i}$ of the (diagonal) twist matrix ${\bf g}$ as on parameters. From commutativity of the ${\sf T}_{\lambda}(x)$ for all $x,\lambda$ it follows that the operators ${\sf T}(x;{\bf t})$ commute for all $x$, ${\bf t}$. In terms of the co-derivatives, the master $T$-operator can be represented in the form
${\sf T}(x;{\bf t})=(x-x_{N}+\eta D_{N})\ldots(x-x_{1}+\eta D_{1})\exp\Bigl{(}\sum_{k\geq 1}t_{k}{\rm tr}\,{\bf g}^{k}\Bigr{)}$ (4.2)
(to see this, one should use (3.24) and the Cauchy-Littlewood identity (3.11)).
Obviously, ${\sf T}(x;0)={\sf T}_{\emptyset}(x)=\phi(x)$. Acting to ${\sf T}(x;{\bf t})$ by differential operators in $t_{k}$ at ${\bf t}=0$, one can reproduce all the transfer matrices ${\sf T}_{\lambda}(x)$. For example,
${\sf T}_{(1)}(x)=\partial_{t_{1}}{\sf T}(x;{\bf t})\Bigr{|}_{{\bf t}=0},\qquad{\sf T}_{(2)}(x)=\frac{1}{2}\left(\partial_{t_{1}}^{2}+\partial_{t_{2}}\right){\sf T}(x;{\bf t})\Bigr{|}_{{\bf t}=0}.$ (4.3)
The general formula is a direct consequence of orthogonality of the Schur functions which can be easily derived from the Cauchy-Littlewood identity in the form
$s_{\lambda}(\tilde{\partial})s_{\mu}({\bf t})\Bigr{|}_{{\bf t}=0}=\delta_{\lambda\mu},$
where ~=$\tilde{\partial}=\{\partial_{t_{1}},\frac{1}{2}\partial_{t_{2}},\frac{1}{3}\partial_{t_{3}},\ldots\}$. Acting by $s_{\lambda}(\tilde{\partial})$ to the both sides of (4.1), we conclude that
${\sf T}_{\lambda}(x)=s_{\lambda}(\tilde{\partial}){\sf T}(x;{\bf t})\Bigr{|}_{{\bf t}=0}.$ (4.4)
Below we use the standard notation
$\begin{array}[]{c}{\bf t}\pm[z^{-1}]=\Bigl{\{}t_{1}\pm z^{-1},t_{2}\pm\frac{1}{2}z^{-2},t_{3}\pm\frac{1}{3}z^{-3},\ldots\Bigr{\}}.\end{array}$ (4.5)
From (4.2) it follows that ${\sf T}(x,0\pm[z^{-1}])$ is the generating series for the transfer matrices corresponding to the diagrams of the form of one row or one column:
${\sf T}(x,[z^{-1}])=\sum_{s=0}^{\infty}z^{-s}{\sf T}_{s}(x),\qquad{\sf T}(x,-[z^{-1}])=\sum_{a=0}^{n}(-z)^{-a}{\sf T}^{a}(x).$ (4.6)
Indeed, we have for the first equation:
$\sum_{s\geq 0}z^{-s}{\sf T}_{s}(x)=\sum_{s\geq 0}z^{-s}h_{s}(\tilde{\partial}){\sf T}(x;{\bf t})\Bigr{|}_{{\bf t}=0}=\exp\Bigl{(}\sum_{k\geq 1}\frac{1}{k}z^{-k}\partial_{t_{k}}\Bigr{)}{\sf T}(x;{\bf t})\Bigr{|}_{{\bf t}=0}={\sf T}(x,[z^{-1}]),$
where we have used the generating function (3.9). The second identity is proved in a similar way.
More generally, from (4.2) we also have:
$\begin{array}[]{l}\displaystyle{{\sf T}(x;{\bf t}-[z^{-1}])=(x\!-\!x_{N}\!+\!\eta D_{N})\ldots(x\!-\!x_{1}\!+\!\eta D_{1})\Bigl{[}\det({\sf I}-z^{-1}{\sf g})\exp\Bigl{(}\sum_{k\geq 1}t_{k}{\rm tr}\,{\bf g}^{k}\Bigr{)}\Bigr{]},}\\ \\ \displaystyle{{\sf T}(x;{\bf t}+[z^{-1}])=(x\!-\!x_{N}\!+\!\eta D_{N})\ldots(x\!-\!x_{1}\!+\!\eta D_{1})\Bigl{[}\frac{1}{\det({\sf I}-z^{-1}{\sf g})}\exp\Bigl{(}\sum_{k\geq 1}t_{k}{\rm tr}\,{\bf g}^{k}\Bigr{)}\Bigr{]}.}\end{array}$ (4.7)
From the right-hand side of the second formula it is clear that ${\sf T}(x;{\bf t}+[z^{-1}])$ has poles if $z$ equals any eigenvalue of ${\bf g}$. Since each co-derivative raises the order of a pole, these poles are multiple.
4.2 Bilinear functional relations for the master $T$-operator
Let us pass to the most important property of the master $T$-operator, which establishes a close connection with the theory of classical integrable nonlinear partial differential equations. Namely, it was proved in [15] that the CBR relations (3.19) (or (3.20)) mean that any eigenvalue of the master $T$-operator as a function of the variables $x$, ${\bf t}$ is a tau-function of the classical modified KP (mKP) hierarchy known in the theory of soliton equations. This follows from the fact, proved in [15] (see also [49]), that if the coefficients $c_{\lambda}(x)$ in the expansion
$\tau(x,{\bf t})=\sum_{\lambda}c_{\lambda}(x)s_{\lambda}({\bf t})$ (4.8)
obey the quantum Jacobi-Trudi relations
$c_{\lambda}(x)=\left(\prod_{k=1}^{\lambda_{1}^{\prime}-1}c_{\emptyset}(x-k\eta)\right)^{-1}\det_{1\leq i,j\leq\lambda_{1}^{\prime}}c_{(\lambda_{i}-i+j)}(x-(j\!-\!1)\eta),$ (4.9)
$c_{\lambda}(x)=\left(\prod_{k=1}^{\lambda_{1}-1}c_{\emptyset}(x+k\eta)\right)^{-1}\det_{1\leq i,j\leq\lambda_{1}}c_{(1^{\lambda_{i}^{\prime}-i+j})}(x+(j\!-\!1)\eta),$ (4.10)
then $\tau(x,{\bf t})$ is a tau-function of the mKP hierarchy, in which $x$ is identified with the “zeroth time”. This fact allows one to regard the master $T$-operator (4.1) as an operator-valued tau-function meaning that any its eigenvalue is a solution to the mKP hierarchy represented in the bilinear form.
Note that there is a freedom to multiply the tau-function by $A^{x}$, where $A$ is any non-zero constant: the transformation $c_{\lambda}(x)\to A^{x}c_{\lambda}(x)$ does not spoil the relations (4.9), (4.10). We call the tau-functions that differ by a factor of the form $A^{x}$ equivalent.
It follows from the bilinear formalism [31, 32] that the generating bilinear relation for the master $T$-operator can be written in the form
[…]
$\xi({\bf t},z)=\sum_{k\geq 1}t_{k}z^{k}.$ (4.12)
The integration contour $C_{\infty}$ is a big circle of radius $R\to\infty$ which separates the singularities coming from the ${\sf T}$-multipliers and the exponential function. For some special values of $x-x^{\prime}$ and ${\bf t}-{\bf t}^{\prime}$, the integral can be calculated by residue calculus and equation (4.11) is thus a source of various bilinear relations of the Hirota-Miwa type. For example, one can put $x^{\prime}=x-\eta$, ${\bf t}^{\prime}={\bf t}-[z_{1}^{-1}]-[z_{2}^{-1}]$, then
$e^{\xi({\bf t}-{\bf t}^{\prime},z)}=\frac{z_{1}z_{2}}{(z_{1}-z)(z_{2}-z)}$
and the residue calculus gives the following 3-term bilinear equation:
$\begin{array}[]{c}z_{2}{\sf T}\left(x+\eta;{\bf t}-[z_{2}^{-1}]\right){\sf T}\left(x;{\bf t}-[z_{1}^{-1}]\right)-z_{1}{\sf T}\left(x+\eta;{\bf t}-[z_{1}^{-1}]\right){\sf T}\left(x;{\bf t}-[z_{2}^{-1}]\right)\\ \\ +\,(z_{1}-z_{2}){\sf T}(x+\eta;{\bf t}){\sf T}\left(x;{\bf t}-[z_{1}^{-1}]-[z_{2}^{-1}]\right)=0.\end{array}$ (4.13)
In another form, this equation for ${\sf T}(x;{\bf t})$ first appeared in [30]. The same equation arises from (4.11) if one puts $x^{\prime}=x+\eta$, ${\bf t}^{\prime}={\bf t}-[z_{1}^{-1}]-[z_{2}^{-1}]$.
We have seen that the behavior of the master $T$-operator ${\sf T}(x;{\bf t})$ as a function of ${\bf t}$ in an infinitesimally small neighborhood of the point ${\bf t}=0$ contains all information about the transfer matrices ${\sf T}_{\lambda}(x)$. Below we will see that analyzing its analytic properties near some other points other than ${\bf t}=0$, one can recover Baxter’s $Q$-operators. To be more precise, it was shown in [15] that ${\sf T}(x;[z_{1}^{-1}]+\ldots+[z_{m}^{-1}])$ regarded as a function of $z_{i}$’s has multiple poles when any of the $z_{i}$’s coincides with any one of the twist parameters $g_{j}$’s and residues at these poles can be identified with the $Q$-operators.
5 The mKP hierarchy
The mKP hierarchy can be defined as a system of evolution equations for the Lax operator ${\bf L}(x)$ which is a pseudo-difference operator of the form
${\bf L}(x)=e^{\eta\partial_{x}}+\sum_{j\geq 0}u_{j}(x)e^{-j\eta\partial_{x}}.$ (5.1)
Here $e^{\pm\eta\partial_{x}}$ is the shift operator acting to functions of $x$ as $e^{\pm\eta\partial_{x}}f(x)=f(x\pm\eta)$. The evolution equations (the Lax equations) in the times $t_{k}$ are as follows:
$\partial_{t_{k}}{\bf L}(x)=[{\bf B}_{k},\,{\bf L}(x)],\qquad{\bf B}_{k}=({\bf L}^{k}(x))_{\geq 0},$ (5.2)
where the notation ${\Bigl{(}\sum_{j}v_{j}e^{j\eta\partial_{x}}\Bigr{)}_{\geq 0}}$ means ${\sum_{j\geq 0}v_{j}e^{j\eta\partial_{x}}}$, so ${\bf B}_{k}$ is a difference operator of order $k$.
An important role in the theory is played by the so-called wave (or dressing) operator ${\bf W}(x)$ which is a pseudo-difference operator of the form
${\bf W}(x)=1+\sum_{j\geq 1}w_{j}(x)e^{-j\eta\partial_{x}}$ (5.3)
such that
${\bf L}(x)={\bf W}(x)e^{\eta\partial_{x}}{\bf W}^{-1}(x).$ (5.4)
This representation is often interpreted as “dressing” of the trivial Lax operator $e^{\eta\partial_{x}}$ by ${\bf W}$. The wave operator obeys the evolution equation
$\partial_{t_{k}}{\bf W}(x)={\bf B}_{k}(x){\bf W}(x)-{\bf W}(x)e^{k\eta\partial_{x}},$ (5.5)
so it depends also on ${\bf t}$: ${\bf W}(x)={\bf W}(x,{\bf t})$. The inverse operator ${\bf W}^{-1}$ is of the form
${\bf W}^{-1}(x)=1+\sum_{j\geq 1}e^{-j\eta\partial_{x}}w_{j}^{*}(x+\eta)=1+\sum_{j\geq 1}w_{j}^{*}(x-(j-1)\eta)e^{-j\eta\partial_{x}}.$ (5.6)
[…]
$\begin{array}[]{l}\psi(x,{\bf t};z)={\bf W}(x)z^{x/\eta}e^{\xi({\bf t},z)},\\ \\ \psi^{*}(x,{\bf t};z)=({\bf W}^{-1}(x-\eta))^{{\dagger}}z^{-x/\eta}e^{-\xi({\bf t},z)}.\end{array}$ (5.7)
The operation $(\ldots)^{{\dagger}}$ is defined as $(f(x)e^{k\eta\partial_{x}})^{{\dagger}}=e^{-k\eta\partial_{x}}f(x)$ and is extended by linearity to all pseudo-difference operators. From (5.7) we see that the wave function and its adjoint have the following expansions as $z\to\infty$:
$\begin{array}[]{l}\displaystyle{\psi(x,{\bf t};z)=z^{x/\eta}e^{\xi({\bf t},z)}\Bigl{(}1+\frac{w_{1}(x,{\bf t})}{z}+\frac{w_{2}(x,{\bf t})}{z^{2}}+\ldots\Bigr{)},}\end{array}$ (5.8)
$\begin{array}[]{l}\displaystyle{\psi^{*}(x,{\bf t};z)=z^{-x/\eta}e^{-\xi({\bf t},z)}\Bigl{(}1+\frac{w_{1}^{*}(x,{\bf t})}{z}+\frac{w_{2}^{*}(x,{\bf t})}{z^{2}}+\ldots\Bigr{)}.}\end{array}$ (5.9)
It can be proved that there exists a function $\tau(x,{\bf t})$ such that the wave function and its adjoint are expressed through it in the following way:
$\psi(x,{\bf t};z)=z^{x/\eta}e^{\xi({\bf t},z)}\frac{\tau(x,{\bf t}-[z^{-1}])}{\tau(x,{\bf t})},$ (5.10)
$\psi^{*}(x,{\bf t};z)=z^{-x/\eta}e^{-\xi({\bf t},z)}\frac{\tau(x,{\bf t}+[z^{-1}])}{\tau(x,{\bf t})},$ (5.11)
where we use the notation introduced in (4.5). This function is called the tau-function. It plays a fundamental role in the theory because equations (5.10), (5.11) allow one to express all the coefficients $w_{j}(x,{\bf t})$ of the wave operator (and thus the coefficients $u_{j}(x,{\bf t})$ of the Lax operator) in terms of it.
The tau-function satisfies the integral bilinear equation of the form (4.11):
[…]
$\begin{array}[]{c}z_{2}\tau\left(x+\eta;{\bf t}-[z_{2}^{-1}]\right)\tau\left(x;{\bf t}-[z_{1}^{-1}]\right)-z_{1}\tau\left(x+\eta;{\bf t}-[z_{1}^{-1}]\right)\tau\left(x;{\bf t}-[z_{2}^{-1}]\right)\\ \\ +\,(z_{1}-z_{2})\tau\Bigl{(}x+\eta;{\bf t}\Bigr{)}\tau\left(x;{\bf t}-[z_{1}^{-1}]-[z_{2}^{-1}]\right)=0.\end{array}$ (5.13)
The wave function and its adjoint satisfy an infinite number of differential-difference equations. The simplest ones are
$\begin{array}[]{l}\phantom{-}\partial_{t_{1}}\psi(x,{\bf t};z)=\psi(x+\eta,{\bf t};z)+v(x,{\bf t})\psi(x,{\bf t};z),\\ \\ -\partial_{t_{1}}\psi^{*}(x,{\bf t};z)=\psi^{*}(x-\eta,{\bf t};z)+v(x-\eta,{\bf t})\psi^{*}(x,{\bf t};z),\end{array}$ (5.14)
where
$v(x,{\bf t})=\partial_{t_{1}}\log\frac{\tau(x+\eta,{\bf t})}{\tau(x,{\bf t})}.$ (5.15)
After the substitutions (5.10), (5.11) both of them become equivalent to the following bilinear equation for the tau-function:
$\begin{array}[]{l}\phantom{a}z\tau\Bigl{(}x+\eta,{\bf t}\Bigr{)}\tau\Bigl{(}x,{\bf t}-[z^{-1}]\Bigr{)}-z\tau\Bigl{(}x,{\bf t}\Bigr{)}\tau\Bigl{(}x+\eta,{\bf t}-[z^{-1}]\Bigr{)}\\ \\ =\tau\Bigl{(}x,{\bf t}-[z^{-1}]\Bigr{)}\partial_{t_{1}}\tau\Bigl{(}x+\eta,{\bf t}\Bigr{)}-\tau\Bigl{(}x+\eta,{\bf t}\Bigr{)}\partial_{t_{1}}\tau\Bigl{(}x,{\bf t}-[z^{-1}]\Bigr{)},\end{array}$ (5.16)
which follows from (5.13) in the limit $z_{2}\to\infty$ if one puts $z_{1}=z$.
The mKP hierarchy has a lot of solutions of very different nature. To specify the class of solutions to which eigenvalues of transfer matrices of quantum spin chains belong, it is important to note that for $GL(n)$-invariant models the sum in the series (4.8) goes over Young diagrams $\lambda$ with not more than $n$ non-empty rows:
$\tau(x,{\bf t})=\sum_{\lambda,\,\lambda_{1}^{\prime}\leq n}c_{\lambda}(x)s_{\lambda}({\bf t}).$ (5.17)
It is easy to see that in this case the series in inverse powers of $z$ for $\tau(x,{\bf t}-[z^{-1}])$ (and thus the series (5.8) for the wave function) truncates at the $n$-th term. (This is also obvious from (4.7).) As far as analytical properties in $z$ of the function $\tau(x,{\bf t}+[z^{-1}])$ are concerned, the second equation in (4.7) suggests that it has multiple poles when $z$ is equal to any eigenvalue of the twist matrix. So, characterizing the relevant class of solutions, we should take into account that:
  • a)
    The series (5.8) for the wave function truncates at the $n$-th term (i.e., the wave function, as a function of $z$, has pole of order $n$ at $z=0$);
  • b)
    The tau-function is a polynomial in $x$ of degree $N$.
  • c)
[…]
In the next section we study solutions of this class in more details.
6 Polynomial solutions to the mKP hierarchy
As it was argued in the previous section, we are interested in solutions to the mKP hierarchy such that the tau-function $\tau(x,{\bf t})$ is a polynomial or quasi-polynomial [^3] By quasi-polynomial in $x$ we mean a polynomial multiplied by $A^{x}$, where $A$ is a constant. in $x$. For the KP equation, such solutions were constructed by Krichever. Here we briefly recall this construction, modifying it for the mKP case.
6.1 The wave function
First of all, we consider the wave functions $\psi(x,{\bf t};z)$ such that the series (5.8) in inverse powers of $z$ truncates at the $n$-th term:
$\psi(x,{\bf t};z)=z^{x/\eta}e^{\xi({\bf t},z)}\Bigl{(}1+\frac{w_{1}(x,{\bf t})}{z}+\ldots+\frac{w_{n}(x,{\bf t})}{z^{n}}\Bigr{)}.$ (6.1)
Let $p_{1},\ldots,p_{n}$ be $n$ distinct points in . (Later they will be identified with eigenvalues $g_{i}$ of the twist matrix ${\bf g}$; the exact relation is $p_{i}=g_{n-i+1}$, $i=1,\ldots,n$.) With each point $p_{i}$ we associate an integer number $M_{i}\geq 0$ and $n\times(M_{i}+1)$ rectangular matrix with matrix elements $a_{im}$ ($i=1,\ldots,n$, $m=0,1,\ldots,M_{i}$) which are parameters of the solution. We assume that $a_{i0}\neq 0$. With these data at hand, we impose the following $n$ conditions to the wave function $\psi(x,{\bf t};z)$:
$\sum_{m=0}^{M_{i}}a_{im}\,\partial_{z}^{m}\psi(x,{\bf t};z)\Bigr{|}_{z=p_{i}}\!\!=0\,,\quad i=1,\ldots,n$ (6.2)
We call the set of parameters $\{p_{i}\}$ and $a_{im}$ the Krichever data of the solution and the conditions (6.2) the Krichever conditions. Note that the parameters $a_{i0}$, if they are non-zero, can always be put equal to 1 without any loss of generality, so in what follows we assume that $a_{i0}=1$.
The set of conditions (6.2) yields a system of $n$ linear equations for $n$ coefficients $w_{k}$ which allows one to fix them as functions of $x,{\bf t}$. From the general theory of the mKP hierarchy it then follows that the tau-function associated with the wave function solves the mKP hierarchy. The coefficients $w_{k}$ turn out to be rational functions of their arguments while the tau-function is a polynomial multiplied by the exponential function of a linear combination of the variables $x,{\bf t}$ (a quasi-polynomial). From the algebro-geometric point of view this solution is associated with a highly singular algebraic curve which is the Riemann sphere with cusp singularities at the points $p_{i}$.
It is easy to see that conditions (6.2) are equivalent to the system of linear equations
$A_{i}(x,{\bf t})+\sum_{k=1}^{n}A_{i}(x-k\eta,{\bf t})w_{k}=0$ (6.3)
for the coefficients $w_{k}$ of the wave function, where
$\left.A_{i}(x,{\bf t})=\sum_{m=0}^{M_{i}}a_{im}\partial_{z}^{m}\left(z^{x/\eta}e^{\xi({\bf t},z)}\right)\right|_{z=p_{i}}.$ (6.4)
As is seen from this representation, each function $A_{i}(x,{\bf t})$ is a polynomial in $x$ of degree $M_{i}$ multiplied by the exponential factor $p_{i}^{x/\eta}$. Note that $A_{i}(0,0)=a_{i0}$. The system (6.3) can be solved by applying the Cramer’s rule. This results in the following determinant representation for the wave function:
$\psi(x,{\bf t};z)=z^{x/\eta}e^{\xi({\bf t},z)}\,\frac{\det\left(\begin{array}[]{cccc}1&z^{-1}&\ldots&z^{-n}\\ A_{1}(x,{\bf t})&A_{1}(x\!-\!\eta,{\bf t})&\ldots&A_{1}(x\!-\!n\eta,{\bf t})\\ \vdots&\vdots&\ddots&\vdots\\ A_{n}(x,{\bf t})&A_{n}(x\!-\!\eta,{\bf t})&\ldots&A_{n}(x\!-\!n\eta,{\bf t})\end{array}\right)}{\det\left(\begin{array}[]{ccc}A_{1}(x-\eta,{\bf t})&\ldots&A_{1}(x-n\eta,{\bf t})\\ \vdots&\ddots&\vdots\\ A_{n}(x-\eta,{\bf t})&\ldots&A_{n}(x-n\eta,{\bf t})\end{array}\right)}.$ (6.5)
From the definition (6.4) we have;
$\begin{array}[]{ll}A_{k}(x,{\bf t}-[z^{-1}])=&\displaystyle{\left.\sum_{m=0}^{M_{k}}a_{km}\partial_{\zeta}^{m}\left(\zeta^{x/\eta}e^{\xi({\bf t},\zeta)}\Bigl{(}1-\frac{\zeta}{z}\Bigr{)}\right)\right|_{\zeta=p_{i}}}\\ &\\ &=\,A_{k}(x,{\bf t})-A_{k}(x+\eta,{\bf t})z^{-1}.\end{array}$ (6.6)
Using this, it is straightforward to verify that equation (6.5) agrees with the general relation (5.10), with the tau-function being given by the determinant in the denominator:
$\tau(x,{\bf t})=\det_{i,j=1,\ldots,n}A_{i}(x\!-\!j\eta,\,{\bf t})=\det\left(\begin{array}[]{ccc}A_{1}(x-\eta,{\bf t})&\ldots&A_{1}(x-n\eta,{\bf t})\\ \vdots&\ddots&\vdots\\ A_{n}(x-\eta,{\bf t})&\ldots&A_{n}(x-n\eta,{\bf t})\end{array}\right).$ (6.7)
It is a polynomial in $x$ of degree $N=\sum\limits_{j=1}^{n}M_{j}$ multiplied by $\prod\limits_{i=1}^{n}p_{i}^{x/\eta}e^{\xi({\bf t},p_{i})}$.
It follows from (6.5) that the last coefficient in (6.1), $w_{n}$, is given by
$w_{n}(x,{\bf t})=(-1)^{N}\frac{\tau(x+\eta,{\bf t})}{\tau(x,{\bf t})}.$ (6.8)
We also note the property
$\partial_{t_{1}}A_{i}(x,{\bf t})=A_{i}(x+\eta,{\bf t})$ (6.9)
from which one can see that the first coefficient in (6.1), $w_{1}$, is given by
$w_{1}(x,{\bf t})=-\partial_{t_{1}}\log\tau(x,{\bf t}).$ (6.10)
6.2 The adjoint wave function
Similarly to (6.6), we have:
$A_{k}(x,{\bf t}+[z^{-1}])=\left.\sum_{m=0}^{M_{k}}a_{km}\partial_{\zeta}^{m}\left(\frac{\zeta^{x/\eta}e^{\xi({\bf t},\zeta)}}{1-\zeta/z}\right)\right|_{\zeta=p_{k}},$ (6.11)
Note that this function regarded as a function of $z$ has a pole of order $M_{k}+1$ at $z=p_{k}$. The principal term is
$A_{k}(x,{\bf t}+[z^{-1}])=\frac{M_{k}!\,a_{kM_{k}}p_{k}^{x/\eta+1}e^{\xi({\bf t},p_{k})}}{(z-p_{k})^{M_{k}+1}}\,+\,\ldots\,.$ (6.12)
We also see from (6.11) that the function $A_{k}(x,{\bf t}+[z^{-1}])$ is a rational function of $z$ with simple zero at $z=0$.
In order to obtain a more detailed information about the pole structure of this function, let us add and subtract the term $z^{x/\eta}e^{\xi({\bf t},z)}$ in the numerator in (6.11) and separate the nonsingular part from the singular one:
$\begin{array}[]{l}\displaystyle{A_{k}(x,{\bf t}+[z^{-1}])=z^{x/\eta+1}e^{\xi({\bf t},z)}\sum_{m=0}^{M_{k}}\frac{m!\,a_{km}}{(z-p_{k})^{m+1}}}\\ \\ \phantom{aaaaaaaaaaaaaaaaaaaa}\displaystyle{\left.-z\Bigl{(}\sum_{m=0}^{M_{k}}a_{km}\partial_{\zeta}^{m}\Bigr{)}\,\frac{z^{x/\eta}e^{\xi({\bf t},z)}\!-\!\zeta^{x/\eta}e^{\xi({\bf t},\zeta)}}{z-\zeta}\,\right|_{\zeta=p_{k}}.}\end{array}$ (6.13)
The first sum gives the multiple pole structure at the point $p_{k}$ while the second term is obviously regular at $p_{k}$ and has possible (essential) singularities and branching only at 00 and $\infty$.
Rewriting (6.6) in the form
$A_{k}(x,{\bf t}+[z^{-1}])=A_{k}(x,{\bf t})+z^{-1}A_{k}(x+\eta,{\bf t}+[z^{-1}]),$
it is straightforward to check that
$\tau(x,{\bf t}+[z^{-1}])=\det\left(\begin{array}[]{cccc}A_{1}\left(x\!-\!\eta,{\bf t}\!+\![z^{-1}]\right)&A_{1}(x\!-\!2\eta,{\bf t})&\ldots&A_{1}(x\!-\!n\eta,{\bf t})\\ A_{2}\left(x\!-\!\eta,{\bf t}\!+\![z^{-1}]\right)&A_{2}(x\!-\!2\eta,{\bf t})&\ldots&A_{2}(x\!-\!n\eta,{\bf t})\\ \vdots&\vdots&\ddots&\vdots\\ A_{n}\left(x\!-\!\eta,{\bf t}\!+\![z^{-1}]\right)&A_{n}(x\!-\!2\eta,{\bf t})&\ldots&A_{n}(x\!-\!n\eta,{\bf t}).\end{array}\right).$ (6.14)
Expanding (6.11) in powers of $z$, we get:
$A_{k}(x,{\bf t}+[z^{-1}])=A_{k}(x,{\bf t})+A_{k}(x+\eta,{\bf t})z^{-1}+A_{k}(x+2\eta,{\bf t})z^{-2}+\ldots\,,$
and so the expansion of $\tau(x,{\bf t}+[z^{-1}])$ around $\infty$ reads
$\tau(x,{\bf t}+[z^{-1}])=\sum_{s=0}^{\infty}z^{-s}\det\left(\begin{array}[]{cccc}A_{1}\left(x\!+(s-1)\eta,{\bf t}\right)&A_{1}(x\!-\!2\eta,{\bf t})&\ldots&A_{1}(x\!-\!n\eta,{\bf t})\\ A_{2}\left(x\!+(s-1)\eta,{\bf t}\right)&A_{2}(x\!-\!2\eta,{\bf t})&\ldots&A_{2}(x\!-\!n\eta,{\bf t})\\ \vdots&\vdots&\ddots&\vdots\\ A_{n}\left(x\!+(s-1)\eta,{\bf t}\right)&A_{n}(x\!-\!2\eta,{\bf t})&\ldots&A_{n}(x\!-\!n\eta,{\bf t})\end{array}\right).$ (6.15)
We thus see that the adjoint wave function has the determinant representation
$\psi^{*}(x,{\bf t};z)=z^{-x/\eta}e^{-\xi({\bf t},z)}\,\frac{\det\left(\begin{array}[]{cccc}A_{1}\left(x\!-\!\eta,{\bf t}\!+\![z^{-1}]\right)&A_{1}(x\!-\!2\eta,{\bf t})&\ldots&A_{1}(x\!-\!n\eta,{\bf t})\\ A_{2}\left(x\!-\!\eta,{\bf t}\!+\![z^{-1}]\right)&A_{2}(x\!-\!2\eta,{\bf t})&\ldots&A_{2}(x\!-\!n\eta,{\bf t})\\ \vdots&\vdots&\ddots&\vdots\\ A_{n}\left(x\!-\!\eta,{\bf t}\!+\![z^{-1}]\right)&A_{n}(x\!-\!2\eta,{\bf t})&\ldots&A_{n}(x\!-\!n\eta,{\bf t})\end{array}\right)}{\det\left(\begin{array}[]{cccc}A_{1}\left(x\!-\!\eta,{\bf t}\right)&A_{1}(x\!-\!2\eta,{\bf t})&\ldots&A_{1}(x\!-\!n\eta,{\bf t})\\ A_{2}\left(x\!-\!\eta,{\bf t}\right)&A_{2}(x\!-\!2\eta,{\bf t})&\ldots&A_{2}(x\!-\!n\eta,{\bf t})\\ \vdots&\vdots&\ddots&\vdots\\ A_{n}\left(x\!-\!\eta,{\bf t}\right)&A_{n}(x\!-\!2\eta,{\bf t})&\ldots&A_{n}(x\!-\!n\eta,{\bf t})\end{array}\right).}$ (6.16)
This function has multiple poles at the points $p_{i}$. Since the function $A_{k}(x,{\bf t}+[z^{-1}])$ has a simple zero at $z=0$, it is clear from (6.7) that the function $\tau(x,{\bf t}+[z^{-1}])$ and thus the function $z^{x/\eta}e^{\xi({\bf t},z)}\psi^{*}(x,{\bf t};z)$ has zero of order $n$ at $z=0$.
Let us introduce the notation
[…]
for the minor of the $n\!\times\!n$ matrix $A_{i}(x+(1-j)\eta)$, $1\leq i,j\leq n$. Then, expanding the determinant in the numerator of (6.16) in the first column, we obtain:
$\psi^{*}(x,{\bf t};z)=\frac{z^{-x/\eta}e^{-\xi({\bf t},z)}}{\tau(x,{\bf t})}\sum_{k=1}^{n}(-1)^{k-1}\hat{A}_{k}(x-2\eta,{\bf t})\,A_{k}(x-\eta,{\bf t}+[z^{-1}]).$
Using (6.13) we can extract the poles:
$\psi^{*}(x,{\bf t};z)=\sum_{k=1}^{n}(-1)^{k-1}\frac{\hat{A}_{k}(x-2\eta,{\bf t})}{\tau(x,{\bf t})}\sum_{m=0}^{M_{k}}\frac{m!\,a_{km}}{(z-p_{k})^{m+1}}\,+\,\mbox{terms regular at all $p_{k}$}.$ (6.18)
Therefore,
$\mathop{\hbox{res}}\limits_{z=p_{k}}\left[(z-p_{k})^{m}\psi^{*}(x,{\bf t};z)\right]=(-1)^{k-1}m!\,a_{km}\,\frac{\hat{A}_{k}(x-2\eta,{\bf t})}{\tau(x,{\bf t})}$ (6.19)
for $m=0,1,\ldots,M_{k}$, or, in terms of the tau-function,
$\mathop{\hbox{res}}\limits_{z=p_{k}}\left[(z-p_{k})^{m}z^{-x/\eta}e^{-\xi({\bf t},z)}\tau(x,{\bf t}+[z^{-1}])\right]=(-1)^{k-1}m!\,a_{km}\,\hat{A}_{k}(x-2\eta,{\bf t}).$ (6.20)
Note that the expression in the right-hand side has the same structure as (6.7) and, therefore, is a (quasi)polynomial tau-function. It is constructed by means of the Krichever data from which the point $p_{k}$ is excluded in the same way as $\tau(x,{\bf t})$. Thus the passage from $\tau(x,{\bf t})$ to $A_{k}(x,{\bf t})$ can be regarded as a Bäcklund transformation.
6.3 Undressing transformations
Our main goal is to show that the nested Bethe ansatz scheme is equivalent to a chain of some special Bäcklund transformations of the initial polynomial mKP solution with the Krichever data $p_{1},\ldots,p_{n}$, $a_{im}$ that “undress” it to the trivial solution by reducing the number of the points $p_{i}$ in succession. Here we present the idea solely in terms of the mKP hierarchy.
Removing a point from the Krichever data of a polynomial solution is a Bäcklund transformation. It sends a (quasi)polynomial tau-function to another one. Moreover, we are going to consider a chain of such transformations. To this end, let us fix some order in the set of points $p_{1},p_{2},\ldots,p_{n}$ and remove first $p_{n}$, then $p_{n-1}$ and so on, up to $p_{1}$. (Finally, removing all the points, we obtain the trivial solution which is a constant.) The key equation that allows one to implement such transformations is (6.20). Let us consider its $m=0$ case and start from removing the point $p_{n}$. Specifically, consider the function
$\tau^{(n-1)}(x,{\bf t})=(-1)^{n-1}\mathop{\hbox{res}}\limits_{z=p_{n}}\left(z^{-x/\eta-1}e^{-\xi({\bf t},z)}\tau(x+\eta,{\bf t}+[z^{-1}])\right).$ (6.21)
According to the $m=0$ case of (6.20), it is equal to
$\tau^{(n-1)}(x,{\bf t})=\det\left(\begin{array}[]{cccc}A_{1}\left(x\!-\!\eta,{\bf t}\right)&A_{1}\,(x\!-\!2\eta,{\bf t})&\ldots&A_{1}(x\!-\!(n-1)\eta,{\bf t})\\ A_{2}\left(x\!-\!\eta,{\bf t}\right)&A_{2}\,(x\!-\!2\eta,{\bf t})&\ldots&A_{2}(x\!-\!(n-1)\eta,{\bf t})\\ \vdots&\vdots&\ddots&\vdots\\ A_{n-1}\left(x\!-\!\eta,{\bf t}\right)&A_{n-1}(x\!-\!2\eta,{\bf t})&\ldots&A_{n-1}\,(x\!-\!(n\!-\!1)\eta,{\bf t})\end{array}\right),$ (6.22)
i.e., to the minor $M_{n,n}$ of the matrix $A_{i}(x+(1-j)\eta)$. Therefore, it is a tau-function, i.e., it satisfies the same bilinear equations of the mKP hierarchy as $\tau(x,{\bf t})$ does and $\tau^{(n)}\rightarrow\tau^{(n-1)}$, where $\tau^{(n)}(x,{\bf t})=\tau(x,{\bf t})$ is the initial member of the undressing chain, is indeed a Bäcklund transformation.
Having at hand $\tau^{(n-1)}(x,{\bf t})$, one can construct the wave function
$\psi^{(n-1)}(x,{\bf t};z)=z^{x/\eta}e^{\xi({\bf t},z)}\,\frac{\tau^{(n-1)}(x,{\bf t}-[z^{-1}])}{\tau^{(n-1)}(x,{\bf t})}$ (6.23)
which obeys the same Krichever conditions (6.2) at the points $p_{1},\ldots,p_{n-1}$ but not at the point $p_{n}$, where no condition is imposed. Note that the wave function $\psi^{(n-1)}$ has the form similar to (6.1) but with a pole at $z=0$ of order $n-1$ rather than $n$, so the number of conditions again matches the number of unknown coefficients. The adjoint wave function is
$\psi^{*(n-1)}(x,{\bf t};z)=z^{-x/\eta}e^{-\xi({\bf t},z)}\,\frac{\tau^{(n-1)}(x,{\bf t}+[z^{-1}])}{\tau^{(n-1)}(x,{\bf t})}.$ (6.24)
It has multiple poles at the points $p_{1},\ldots,p_{n-1}$.
This process can be continued by taking the residue of $z^{-x/\eta-1}e^{-\xi({\bf t},z)}\tau^{(n-1)}(x+\eta,{\bf t}+[z^{-1}])$ at the next point $p_{n-1}$ and introducing the function
$\begin{array}[]{c}\tau^{(n-2)}(x,{\bf t})=(-1)^{n-2}\mathop{\hbox{res}}\limits_{z=p_{n-1}}\left(z^{-x/\eta-1}e^{-\xi({\bf t},z)}\tau^{(n-1)}(x+\eta,{\bf t}+[z^{-1}])\right)\\ \\ \displaystyle{=\mathop{\hbox{res}}\limits_{z_{n-1}=p_{n-1}\atop z_{n}=p_{n}}\Bigl{(}(z_{n-1}z_{n})^{-x/\eta-2}e^{-\xi({\bf t},z_{n-1})-\xi({\bf t},z_{n})}(z_{n}-z_{n-1})\tau^{(n)}(x+2\eta,{\bf t}+[z_{n-1}^{-1}]+[z_{n}^{-1}])\Bigr{)}}\\ \\ =\,\det\limits_{r,s=1,\ldots,n-2}\Bigl{[}A_{r}(x-s\eta,{\bf t})\Bigr{]}\end{array}$ (6.25)
We thus obtain a chain of Bäcklund transformations
$\tau=\tau^{(n)}\rightarrow\tau^{(n-1)}\rightarrow\tau^{(n-2)}\rightarrow\tau^{(n-3)}\rightarrow\ldots\rightarrow\tau^{(1)}\rightarrow\tau^{(0)}=1$ (6.26)
which “undress” the initial solution up to the trivial one. The general recursive formula at the $m$-th level is the same as (6.21) with the change $n\to m$. Solving the recursion relation, we find:
$\begin{array}[]{l}\displaystyle{\tau^{(m)}(x,{\bf t})=\mathop{\hbox{res}}\limits_{z_{j}=p_{j}\atop j=m+1,\ldots,n}\left[\prod_{\alpha=m+1}^{n}(z_{\alpha}^{-x/\eta+m-n}e^{\xi({\bf t},z_{\alpha})})\Delta(z_{m+1},\ldots,z_{n})\right.}\\ \\ \phantom{aaaaaaaaaaaaaaa}\displaystyle{\left.\times\,\tau\Bigl{(}x+(n-m)\eta,{\bf t}+\!\!\sum_{\alpha=m+1}^{n}[z_{\alpha}^{-1}]\Bigr{)}\right],}\end{array}$ (6.27)
where ${\Delta(x_{1},\ldots,x_{m})=\prod\limits_{i>j}^{m}(x_{i}-x_{j})}$ is the Vandermonde determinant. In particular, on the previous to the last level we have
$\tau^{(1)}(x,{\bf t})=A_{1}(x-\eta,{\bf t})$ (6.28)
and on the last level $\tau^{(0)}({\bf t})=1$. The function $\tau^{(m)}(x,{\bf t})$ is a quasipolynomial in $x$ of degree $N_{m}=M_{1}+\ldots+M_{m}$.
Note that equation (6.20) allows one to make the same Bäcklund transformations by picking the coefficients in front of higher order poles of the function $\tau(x,{\bf t}+[z^{-1}])$ at $z=p_{1},\ldots,p_{n}$. The results differ from (6.27) by normalization factors independent of $x,{\bf t}$.
6.4 The dressing chain
In this section we will follow the arrows of the “undressing chain” (6.26) in the inverse direction, from right to left. Then each step is naturally referred to as a “dressing” transformation, which allows one to construct more complicated tau-functions from simpler ones.
The wave function $\psi^{(m)}(x,{\bf t};z)$ at the $m$-th step of the chain is given by (6.23) with the change $n\to m$. We claim that this wave function can be obtained from the previous one, $\psi^{(m-1)}(x,{\bf t};z)$, by action of a first order difference operator. Namely, we have:
$\psi^{(m)}(x,{\bf t};z)=\left(1-\frac{\tau^{(m)}(x+\eta)\tau^{(m-1)}(x-\eta)}{\tau^{(m)}(x+\eta)\tau^{(m-1)}(x)}\,e^{-\eta\partial_{x}}\right)\psi^{(m-1)}(x,{\bf t};z),$ (6.29)
where the dependence of the tau-functions on ${\bf t}$ is not shown explicitly. To prove this equality, we write down the right-hand side,
$z^{x/\eta}e^{\xi({\bf t},z)}\left[\frac{1}{\tau^{(m-1)}(x)}\,\det\left(\begin{array}[]{cccc}1&z^{-1}&\ldots&z^{-(m-1)}\\ A_{1}(x)&A_{1}(x\!-\!\eta)&\ldots&A_{1}(x\!-\!(m-1)\eta)\\ \vdots&\vdots&\ddots&\vdots\\ A_{m-1}(x)&A_{m-1}(x\!-\!\eta)&\ldots&A_{m-1}(u\!-\!n\eta)\end{array}\right)\right.$
$\left.-\,\frac{\tau^{(m)}(x+\eta)}{\tau^{(m)}(x)\tau^{(m-1)}(x)}\det\left(\begin{array}[]{cccc}z^{-1}&z^{-2}&\ldots&z^{-m}\\ A_{1}(x-\eta)&A_{1}(x\!-2\!\eta)&\ldots&A_{1}(x\!-\!m\eta)\\ \vdots&\vdots&\ddots&\vdots\\ A_{m-1}(x-\eta)&A_{m-1}(x\!-2\!\eta)&\ldots&A_{m-1}(x\!-\!m\eta)\end{array}\right)\right],$
and extract the coefficient in front of $z^{-k}$. It is equal to
$w^{(m)}_{k}=(-1)^{k}\left[\frac{\tau^{(m-1),k}(x)}{\tau^{(m-1)}(x)}+\frac{\tau^{(m)}(x+\eta)\tau^{(m-1),k-1}(x-\eta)}{\tau^{(m)}(x)\tau^{(m-1)}(x)}\right],$
where
[…]
To show that
$w_{k}^{(m)}=(-1)^{k}\frac{\tau^{(m),k}(x)}{\tau^{(m)}(x)},$ (6.30)
as it should be, we need the determinant identity
$D[j_{1}j_{2}]D[j_{3}j_{4}]+D[j_{1}j_{4}]D[j_{2}j_{3}]=D[j_{1}j_{3}]D[j_{2}j_{4}]$ (6.31)
valid for determinants $D[ij]$ of square matrices obtained from any $m\times(m+2)$ rectangular matrix $M_{ab}$ by removing its $i$-th and $j$-th columns. This identity is one of the Plücker relations. Let us take the matrix $M_{ab}$ to be $M_{i1}=\delta_{m,i}$, $M_{ij}=A_{i}(x-(j-2)\eta)$ with $i=1,\ldots,m$, $j=2,\ldots,m+2$, and $j_{1}=1$, $j_{2}=2$, $j_{3}=k+1$, $j_{4}=m$. With this choice, the identity (6.31) reads
$\tau^{(m-1),k}(x)\tau^{(m)}(x)+\tau^{(m-1),k-1}(x-\eta)\tau^{(m)}(x+\eta)=\tau^{(m-1)}(x)\tau^{(m),k}(x)$
and thus leads to (6.30).
Regarding (6.29) as a recurrence relation for the wave functions $\psi^{(m)}$, one can factorize the wave operator ${\bf W}^{(n)}$ by representing it as a product of $n$ first order difference operators:
$\begin{array}[]{c}\displaystyle{{\bf W}^{(n)}(x)=\sum_{k=0}^{n}(-1)^{k}\frac{\tau^{(n),k}(x)}{\tau^{(n)}(x)}\,e^{-k\eta\partial_{x}}}\\ \\ =\,(1-U_{n}(x)e^{-\eta\partial_{x}})(1-U_{n-1}(x)e^{-\eta\partial_{x}})\ldots(1-U_{1}(x)e^{-\eta\partial_{x}}),\end{array}$ (6.32)
where
$U_{m}(x)=\frac{\tau^{(m)}(x+\eta)\tau^{(m-1)}(x-\eta)}{\tau^{(m)}(x)\tau^{(m-1)}(x)}.$ (6.33)
In particular,
$\frac{\tau^{(n),1}(x)}{\tau^{(n)}(x)}=\sum_{m=0}^{n}U_{m}(x)=\sum_{m=0}^{n}\frac{\tau^{(m)}(x+\eta)\tau^{(m-1)}(x-\eta)}{\tau^{(m)}(x)\,\tau^{(m-1)}(x)}.$ (6.34)
From the form of the last multiplier in right-hand side of (6.32) it is obvious that $\Psi(x)=\tau^{(1)}(x+\eta,{\bf t})=A_{1}(x,{\bf t})$ is a solution to the difference equation
${\bf W}^{(n)}(x)\Psi(x)=0,$ (6.35)
i.e., it holds
$\sum_{k=0}^{n}(-1)^{k}\frac{\tau^{(n),k}(x)}{\tau^{(n)}(x)}\,\tau^{(1)}(x-(k-1)\eta,{\bf t})=0.$ (6.36)
The representation of equation (6.35) in the form
$\det\left(\begin{array}[]{cccc}\Psi(x)&\Psi(x-\eta)&\ldots&\Psi(x-n\eta)\\ A_{1}(x)&A_{1}(x\!-\!\eta)&\ldots&A_{1}(x\!-\!n\eta)\\ \vdots&\vdots&\ddots&\vdots\\ A_{n}(x)&A_{n-1}(x\!-\!\eta)&\ldots&A_{n}(x\!-\!n\eta)\end{array}\right)=0$ (6.37)
makes it obvious that the other $n-1$ solutions are $A_{2}(x,{\bf t}),\,A_{3}(x,{\bf t}),\ldots,A_{n}(x,{\bf t})$. (However, this is not so easy to see from (6.32).)
Let us obtain the relation connecting tau-functions at two neighboring levels of the dressing chain. The easiest way to do this is to substitute the expression for the wave function
$\psi^{(m)}(x,{\bf t};z)=z^{x/\eta}e^{\xi({\bf t},z)}\,\frac{\tau^{(m)}(x,{\bf t}-[z^{-1}])}{\tau^{(m)}(x,{\bf t})}$
into (6.29). This gives the following bilinear equation:
$\begin{array}[]{l}\tau^{(m-1)}(x,{\bf t}-[z^{-1}])\tau^{(m)}(x,{\bf t})-z^{-1}\tau^{(m-1)}(x-\eta,{\bf t}-[z^{-1}])\tau^{(m)}(x+\eta,{\bf t})\\ \\ \phantom{aaaaaaaaaaaaaaaaaaaaa}=\tau^{(m-1)}(x,{\bf t})\tau^{(m)}(x,{\bf t}-[z^{-1}]).\end{array}$ (6.38)
It suggests that the variable $m$ can be interpreted as a discrete “time” corresponding to the Bäcklund flow. In (6.38), $m$ takes the finite set of integer values $1,2,\ldots,n$. It is sometimes convenient to extend this set by including the values $m=0$ and $m=n+1$; in this case one should put $\tau^{(-1)}(x,{\bf t})=\tau^{(n+1)}(x,{\bf t})=0$.
Let us denote $\tau^{(m)}(x,{\bf t}-[z^{-1}])=\rho^{(m)}(x)$, then equation (6.38) acquires the form
$\rho^{(m-1)}(x)\tau^{(m)}(x)-z^{-1}\rho^{(m-1)}(x-\eta)\tau^{(m)}(x+\eta)=\rho^{(m)}(x)\tau^{(m-1)}(x).$ (6.39)
We recall that $\tau^{(m)}(x)$ are quasipolynomials in $x$ of degree $N_{m}$. Let $v_{\alpha}^{(m)}$, $\alpha=1,\ldots,N_{m}$ be roots of these quasipolynomials, then we can write
$\tau^{(m)}(x)=(p_{1}\ldots p_{m})^{x/\eta}\prod_{\alpha=1}^{N_{m}}(x-v_{\alpha}^{(m)}).$ (6.40)
Putting $x$ equal to $v_{\alpha}^{(m)}$, $v_{\alpha}^{(m)}-\eta$, $v_{\alpha}^{(m-1)}$ in (6.39), so that only two of the three terms survive, we get the system of equations
$\left\{\begin{array}[]{l}-z^{-1}\rho^{(m-1)}(v_{\alpha}^{(m)}-\eta)\tau^{(m)}(v_{\alpha}^{(m)}+\eta)=\rho^{(m)}(v_{\alpha}^{(m)})\tau^{(m-1)}(v_{\alpha}^{(m)}),\\ \\ \rho^{(m-1)}(v_{\alpha}^{(m)}-\eta)\tau^{(m)}(v_{\alpha}^{(m)}-\eta)=\rho^{(m)}(v_{\alpha}^{(m)}-\eta)\tau^{(m-1)}(v_{\alpha}^{(m)}-\eta),\\ \\ \rho^{(m-1)}(v_{\alpha}^{(m-1)})\tau^{(m)}(v_{\alpha}^{(m-1)})=z^{-1}\rho^{(m-1)}(v_{\alpha}^{(m-1)}-\eta)\tau^{(m)}(v_{\alpha}^{(m-1)}+\eta).\end{array}\right.$
Dividing the first equation by the second one and using the third equation with the shift $m\to m+1$, one can exclude $\rho$. This yields the following equations for zeros of the tau-functions:
$\frac{\tau^{(m+1)}(v_{\alpha}^{(m)}+\eta)\tau^{(m)}(v_{\alpha}^{(m)}-\eta)\tau^{(m-1)}(v_{\alpha}^{(m)})}{\tau^{(m+1)}(v_{\alpha}^{(m)})\tau^{(m)}(v_{\alpha}^{(m)}+\eta)\tau^{(m-1)}(v_{\alpha}^{(m)}-\eta)}=-1,\quad\alpha=1,\ldots,N_{m},$ (6.41)
or, plugging here (6.40),
$\prod_{\beta=1}^{N_{m+1}}\frac{v_{\alpha}^{(m)}-v_{\beta}^{(m+1)}+\eta}{v_{\alpha}^{(m)}-v_{\beta}^{(m+1)}}\,\prod_{\beta=1,\neq\alpha}^{N_{m}}\frac{v_{\alpha}^{(m)}-v_{\beta}^{(m)}-\eta}{v_{\alpha}^{(m)}-v_{\beta}^{(m)}+\eta}\,\prod_{\beta=1}^{N_{m-1}}\frac{v_{\alpha}^{(m)}-v_{\beta}^{(m-1)}}{v_{\alpha}^{(m)}-v_{\beta}^{(m-1)}-\eta}=\frac{p_{m}}{p_{m+1}},$ (6.42)
in which we recognize the system of nested Bethe equations [^4] We postpone a detailed comparison with the Bethe equations (2.12) obtained via nested Bethe ansatz till the next section.. At $m=1$ the last product in the left-hand side should be put equal to 1.
Note that if equations (6.41) hold, all coefficients $\tau^{(n),k}(x,{\bf t})$ in (6.32) are (quasi)polynomials in $x$. Equivalently, equations (6.41) could be derived from the requirement that $\tau^{(n),k}(x,{\bf t})$ are quasipolynomials (i.e, that possible poles at the roots of $\tau^{(m)}(x,{\bf t})$ for $m=1,\ldots,n-1$ cancel).
Remarkably, equations (6.41) have the same form as equations of motion for the integrable time discretization of the Ruijsenaars-Schneider model suggested in [50]. In this interpretation, the Bethe roots $v_{\alpha}^{(m)}$ are coordinates of the particles at the $m$-th step of discrete time. In [2] it was shown that the same equations can be obtained from dynamics of zeros of polynomial tau-functions subject to the fully difference version of the Hirota bilinear equation (also known as a fully discrete KP equation). However, there are two important differences between equations (6.41) and the equations obtained in [50]. First, in the former the discrete time $m$ takes only a finite number of values $m=1,\ldots,n-1$. Second, the number of particles $N_{m}$ at the $m$-th time step depends on $m$.
7 Diagonalization of transfer matrices as a chain of Bäcklund transformations
In this section we identify the objects which appeared in the previous section in the context of the mKP hierarchy with the standard objects of the theory of quantum spin chains and rewrite the key formulas of the previous section in the notation adopted in the theory of spin chains.
We denote eigenvalues of the master $T$-operator ${\sf T}(x,{\bf t})$ by $T(x,{\bf t})$. Similarly, let $T_{\lambda}(x)$ be eigenvalues of ${\sf T}_{\lambda}(x)$ (in particular, $T^{a}(x)$ and $T_{s}(x)$ are eigenvalues of ${\sf T}^{a}(x)$ and ${\sf T}_{s}(x)$ respectively). The eigenvalues $T(x,{\bf t})$ are connected with the quasipolynomial tau-functions constructed in the previous section as
$T(x,{\bf t})=(\det{\bf g})^{-x/\eta}\,\tau(x,{\bf t})$ (7.1)
(the factor $(\det{\bf g})^{-x/\eta}$ is necessary to make the left-hand side polynomial in $x$ rather than quasipolynomial). Different eigenvalues correspond to different polynomial solutions of the mKP hierarchy subject to the necessary conditions.
The standard objects of the theory of quantum spin chains can be obtained from the construction of the previous section at ${\bf t}=0$. For example,
$T(x,0)=\phi(x)=(\det{\bf g})^{-x/\eta}\tau(x,0),$
where $\phi(x)$ is the fixed given polynomial (2.5) whose roots are inhomogeneity parameters of the spin chain. We also note that
$T^{a}(x)=(\det{\bf g})^{-x/\eta}\tau^{(n),a}(x,0)$
with $\tau^{(n),a}(x,{\bf t})$ from (6.32).
The wave operator ${\bf W}(x,0)$ in the theory of spin chains is known under the name of the non-commutative generating function for $T^{a}(x)$:
${\bf W}(x,0)=\sum_{a=0}^{n}(-1)^{a}\frac{T^{a}(x)}{\phi(x)}\,e^{-a\eta\partial_{x}}.$ (7.2)
From (4.6) it follows that ${\bf W}^{-1}(x,0)$ is the non-commutative generating series for $T_{s}(x)$:
${\bf W}^{-1}(x,0)=\sum_{s=0}^{\infty}e^{-s\eta\partial_{x}}\,\frac{T_{s}(x+\eta)}{\phi(x+\eta)}=\sum_{s=0}^{\infty}\frac{T_{s}(x-(s-1)\eta)}{\phi(x-(s-1)\eta)}\,e^{-s\eta\partial_{x}}.$ (7.3)
The functions $\tau^{(m)}(x,0)$ from intermediate levels of the dressing chain can be identified with eigenvalues of the Baxter’s $Q$-operators which are quasipolynomials with roots $v_{\alpha}^{(m)}$:
$\tau^{(m)}(x+\eta,0)=Q_{m}(x)=(p_{1}\ldots p_{m})^{x/\eta}\prod_{\alpha=1}^{N_{m}}(x-v_{\alpha}^{(m)}),\quad m=1,\ldots,n-1,$ (7.4)
where $N_{m}=M_{1}+\ldots+M_{m}$. The last one, $Q_{n}(x)$, is the fixed quasipolynomial
$Q_{n}(x)=(\det{\bf g})^{x/\eta}\phi(x).$
The factorization (6.32) of the wave operator at ${\bf t}=0$ in terms of the $Q_{m}$’s acquires the form
$\begin{array}[]{c}\displaystyle{{\bf W}(x,0)=\left(1\!-\!\frac{Q_{n}(x\!+\!\eta)Q_{n-1}(x\!-\!\eta)}{Q_{n}(x)Q_{n-1}(x)}e^{-\eta\partial_{x}}\right)\left(1\!-\!\frac{Q_{n-1}(x\!+\!\eta)Q_{n-2}(x\!-\!\eta)}{Q_{n-1}(x)Q_{n-2}(x)}e^{-\eta\partial_{x}}\right)\ldots}\\ \\ \displaystyle{\ldots\times\left(1-\frac{Q_{2}(x+\eta)Q_{1}(x-\eta)}{Q_{2}(x)Q_{1}(x)}e^{-\eta\partial_{x}}\right)\left(1-\frac{Q_{1}(x+\eta)}{Q_{1}(x)}e^{-\eta\partial_{x}}\right).}\end{array}$ (7.5)
The roots of the $Q_{m}$’s satisfy the nested Bethe equations (6.42) which in fact follow from (7.5) after imposing the condition that the right-hand side is regular when $x$ is equal to any root of $Q_{1}(x),\ldots,Q_{n-1}(x)$. In terms of the $Q_{m}$’s the Bethe equations can be written in the form
$\frac{Q_{m+1}(v_{\alpha}^{(m)}+\eta)Q_{m}(v_{\alpha}^{(m)}-\eta)Q_{m-1}(v_{\alpha}^{(m)})}{Q_{m+1}(v_{\alpha}^{(m)})Q_{m}(v_{\alpha}^{(m)}+\eta)Q_{m-1}(v_{\alpha}^{(m)}-\eta)}=-1,\quad\alpha=1,\ldots,N_{m}.$ (7.6)
In the theory of integrable spin chains, equation (7.5) plays a key role: it allows one to express $T^{a}(x)$ through the eigenvalues of the $Q$-operators (which can be regarded as known quantities as soon as their roots are found from the Bethe equations). With the $T^{a}$’s at hand, eigenvalues of all other transfer matrices ${\sf T}_{\lambda}(x)$ can then be found with the help of the CBR determinant formulas (3.20).
Factorization of the conjugated operator,
${\bf W}^{{\dagger}}(x,0)=\sum_{a=0}^{n}(-1)^{a}\frac{T^{a}(x+a\eta)}{\phi(x+a\eta)}\,e^{a\eta\partial_{x}},$ (7.7)
immediately follows from (7.5):
$\begin{array}[]{c}\displaystyle{{\bf W}^{{\dagger}}(x-\eta,0)=\left(1-\frac{Q_{1}(x+\eta)}{Q_{1}(x)}e^{\eta\partial_{x}}\right)\left(1-\frac{Q_{2}(x+\eta)Q_{1}(x-\eta)}{Q_{2}(x)Q_{1}(x)}e^{\eta\partial_{x}}\right)\ldots}\\ \\ \displaystyle{\ldots\times\left(1\!-\!\frac{Q_{n-1}(x\!+\!\eta)Q_{n-2}(x\!-\!\eta)}{Q_{n-1}(x)Q_{n-2}(x)}e^{\eta\partial_{x}}\right)\left(1\!-\!\frac{Q_{n}(x\!+\!\eta)Q_{n-1}(x\!-\!\eta)}{Q_{n}(x)Q_{n-1}(x)}e^{\eta\partial_{x}}\right).}\end{array}$ (7.8)
Equation (6.36) at ${\bf t}=0$ is the difference equation for $Q_{1}(x)$:
$\sum_{k=0}^{n}(-1)^{k}T^{k}(x)Q_{1}(x-(k-1)\eta)=0.$ (7.9)
The fact that $Q_{1}(x)$ satisfies this equation is also obvious from the form of the last operator multiplier in (7.5). In a similar way, looking at the last multiplier in (7.8), we conclude that $Q_{n-1}(x)$ satisfies the difference equation
$\sum_{a=0}^{n}(-1)^{a}\frac{T^{a}(x+(a-1)\eta)}{\phi(x+(a-1)\eta)}\,\frac{Q_{n-1}(x+(a-1)\eta)}{\phi(x+a\eta)}=0.$ (7.10)
Equations (7.9) and (7.10) generalize the famous Baxter’s $TQ$-relation to $GL(n)$-invariant models with $n>2$. Difference equations for the functions $Q_{i}(x)$ with $22\leq i\leq n-$ 2 also exist but have a more complicated form (see [2]).
As it was already mentioned, the fact that $\Psi(x)=Q_{1}(x+\eta)=A_{1}(x)$ is one of solutions to equation ${\bf W}(x,0)\Psi(x)=0$ is easily seen from the factorization (7.5). The other solutions of this equation are $A_{2}(x),\ldots,A_{n}(x)$, although this is not so easy to see from (7.5). Let us show this for $A_{2}(x)$. To this end, we should change the order of the points $p_{i}$:
$\{p_{1},p_{2},\ldots,p_{n}\}\longrightarrow\{p_{2},p_{1},\ldots,p_{n}\}.$
We denote the $Q$-functions for this order as $\check{Q}_{i}(x)$: $\check{Q}_{1}(x)=A_{2}(x-\eta)$, $\check{Q}_{i}(x)=-Q_{i}(x)$ for $i=2,\ldots,n-1$. Consider the product of the last two operator multipliers in (7.5). It is easy to check that
$\begin{array}[]{c}\displaystyle{\left(1-\frac{Q_{2}(x+\eta)Q_{1}(x-\eta)}{Q_{2}(x)Q_{1}(x)}e^{-\eta\partial_{x}}\right)\left(1-\frac{Q_{1}(x+\eta)}{Q_{1}(x)}e^{-\eta\partial_{x}}\right)}\\ \\ \displaystyle{=\left(1-\frac{Q_{2}(x+\eta)\check{Q}_{1}(x-\eta)}{Q_{2}(x)\check{Q}_{1}(x)}e^{-\eta\partial_{x}}\right)\left(1-\frac{\check{Q}_{1}(x+\eta)}{\check{Q}_{1}(x)}e^{-\eta\partial_{x}}\right).}\end{array}$
From the right-hand side we see that the operator (7.5) indeed kills the function $A_{2}(x)$. For the other solutions the idea of the proof is similar: to change the order of the points $p_{i}$ appropriately, introduce the corresponding functions $\check{Q}_{i}(x)$ and represent the product of several last factors in (7.5) in such a way that the rightmost one would be ${1-(\check{Q}_{1}(x+\eta)/\check{Q}_{1}(x))e^{-\eta\partial_{x}}}$. We omit technical details which can be found in [4].
Finally, let us compare the system of Bethe equations (6.42) obtained as a discrete dynamical system for zeros of tau-functions in the context of the mKP hierarchy with the system (2.12) obtained by the standard methods of the theory of integrable spin chains (the nested Bethe ansatz). Although they look quite similarly, there are differences. In fact the differences are due to the two different (but equivalent) approaches to the problem and can be eliminated merely by a change of notation. Indeed, the standard view on the nested Bethe ansatz is a gradual “undressing” of the original $GL(n)$-problem which is done in $n-1$ steps by transition from $GL(n)$ to $GL(n-1)$, then from $GL(n-1)$ to $GL(n-2)$ and so on, up to the $GL(1)$-model which is trivial. In the mKP-picture this corresponds to the undressing chain (6.26). However, in the mKP context, it is more natural to invert the arrows and follow this chain in the opposite direction, from right to left, then each step is a “dressing” transformation and the chain becomes the dressing chain discussed in Section 6.4. Note that now nothing prevents to continue it infinitely to the left. Therefore, to identify (6.42) and (2.12), we should change the notation in accordance with this understanding. Namely, after the identification
$N_{m}={\cal N}_{n-m},\quad v_{\alpha}^{(m)}=w_{\alpha}^{(n-m)},\quad p_{m}=g_{n-m+1}$ (7.11)
equations (6.42) coincide with (2.12).
8 Connection with the classical Ruijsenaars-Schneider model
We have seen that any eigenvalue $T(x;{\bf t})$ of the master $T$-operator as a function of the times ${\bf t}$ and $t_{0}=x$ is a solution of the mKP hierarchy in the bilinear form (the tau-function), and this tau-function is a (quasi)polynomial in $x$:
$T(x;{\bf t})=e^{t_{1}{\rm tr}\,{\bf g}+t_{2}{\rm tr}\,{\bf g}^{2}+\ldots}\prod_{k=1}^{N}\left(x-x_{k}({\bf t})\right)$ (8.1)
(the exponential factor is restored from the limit $x\to\infty$). The roots of this polynomial depend on $t_{i}$.
The connection with the classical Ruijsenaars-Schneider system of particles becomes clear if one addresses dynamics of zeros of $T(x;{\bf t})$ as functions of the times. The dynamics of zeros of polynomial tau-functions is a well known subject in the theory of integrable nonlinear partial differential equations. In the works by Krichever and others (see [33][40]) it was found that this dynamics is described by equations of motion of integrable many-body systems of the Calogero-Moser and Ruijsenaars-Schneider type. In particular, the dynamics of zeros of the tau-function of the mKP hierarchy of the form (8.1) in the time $t_{k}$ coincides with the dynamics of the Ruijsenaars-Schneider system of particles [41] (which is also known as a relativistic deformation of the Calogero-Moser system [42, 43]) with respect to the $k$-th Hamiltonian flow. For example, the equations of motion in the time $t_{1}$ have the form
$\ddot{x}_{i}=-\sum_{k\neq i}\frac{2\eta^{2}\dot{x}_{i}\dot{x}_{k}}{(x_{i}-x_{k})((x_{i}-x_{k})^{2}-\eta^{2})},$ (8.2)
where the dot means the $t_{1}$-derivative. The parameter $\eta$ has the meaning of the inverse velocity of light. In the limit $\eta\to 0$ one reproduces the Calogero-Moser system of particles.
The Ruijsenaars-Schneider system is Hamiltonian with the Hamiltonian function
${\cal H}=\sum_{i=1}^{N}e^{P_{i}}\prod_{k\neq i}\frac{x_{i}-x_{k}+\eta}{x_{i}-x_{k}}\,,\qquad\{P_{i},x_{k}\}=\delta_{ik}.$
Note that the velocities of the particles are
$\dot{x}_{i}=\frac{\partial{\cal H}}{\partial P_{i}}=e^{P_{i}}\prod_{j\neq i}\frac{x_{i}-x_{j}+\eta}{x_{i}-x_{j}},$ (8.3)
so ${{\cal H}=\sum_{i=1}^{N}\dot{x}_{i}.}$
The system is known to be integrable: there are $N$ independent conserved quantities in involution ${\cal I}_{k}$, $k=1,\ldots,N$, and ${\cal I}_{1}={\cal H}$. Explicitly, they are given by the formula
${\cal I}_{k}=\sum_{I\subset\{1,\ldots,N\},\,|I|=k}\exp\Bigl{(}\sum_{i\in I}P_{i}\Bigr{)}\prod_{i\in I,j\notin I}\frac{x_{i}-x_{j}+\eta}{x_{i}-x_{j}},\quad k=1,\ldots,N.$ (8.4)
In terms of the velocities, the integrals of motion read
${\cal I}_{k}=\sum_{I\subset\{1,\ldots,N\},\,|I|=k}\,\,\Bigl{(}\prod_{i\in I}\dot{x}_{i}\Bigr{)}\prod_{i<j,\,i,j\in I}\frac{(x_{i}-x_{j})^{2}}{(x_{i}-x_{j})^{2}-\eta^{2}}.$ (8.5)
We should recall that the classical $N$-body Ruijsenaars-Schneider model admits a commutation representation in the form of the matrix Lax equation
$\dot{L}=[L,M]$ (8.6)
for $N\!\times\!N$ matrices $L,M$ whose matrix elements are functions of $x_{j}$ and $\dot{x}_{j}$. The matrix $L$ is called the Lax matrix, its explicit form is
$L_{ij}=L_{ij}\Bigl{(}\{\dot{x}_{l}\}_{N},\{x_{l}\}_{N}\Bigr{)}=\frac{\dot{x}_{i}}{x_{i}-x_{j}-\eta},\qquad i,j=1,\ldots,N.$ (8.7)
For our purposes we do not need the explicit form of the matrix $M$. Equations of motion (8.2) are equivalent to the matrix equation (8.6). The Lax equation implies that the time evolution of the Lax matrix $L(0)\rightarrow L(t)$ is an isospectral transformation, i.e., eigenvalues of the Lax matrix (and all symmetric functions of them) are integrals of motion. It is not difficult to see that the characteristic polynomial of the Lax matrix is the generating function of the integrals of motion ${\cal I}_{k}$:
$\det_{N\times N}(zI-L)=z^{N}+\sum_{k=1}^{N}\eta^{-k}\,{\cal I}_{k}z^{N-k},$ (8.8)
where $I$ is the unity matrix. Therefore, the integrals of motion ${\cal I}_{k}$ are given by elementary symmetric polynomials $e_{k}(\xi_{1},\ldots,\xi_{N})$ (see (3.8)) of eigenvalues $\xi_{i}$ of the Lax matrix:
${\cal I}_{k}=(-1)^{k}\eta^{k}e_{k}(\xi_{1},\ldots,\xi_{N}).$ (8.9)
The first formula in (4.3) tells us that the eigenvalue $T(x)$ of the transfer matrix ${\bf T}(x)$ (2.7) is
$T(x)=\partial_{t_{1}}\log T(x;{\bf t})\Bigr{|}_{{\bf t}=0}.$
Plugging here (8.1) and comparing with (2.8), we obtain the following important relation between eigenvalues of the quantum Hamiltonians ${\bf H}_{i}$ (2.9) and initial velocities of the Ruijsenaars-Schneider particles:
$\eta H_{i}=-\dot{x}_{i}(0).$ (8.10)
This relation allows one to make the connection with the Ruijsenaars-Schneider system more precise. For the latter system the standard problem of classical mechanics is to determine time evolution of the $x_{i}$’s from given initial conditions $x_{i}(0),\dot{x}_{i}(0)$, which determine values of all integrals of motion and can be arbitrary. However, in applications to quantum spin chains the problem is posed in a different way. To understand this, recall that in the Krichever’s method which was used in Section 6 for construction of polynomial solutions to the mKP hierarchy only initial coordinates $x_{i}(0)$ (zeros of the tau-function) enter the game as arbitrary parameters (they are implicitly determined from the $a_{im}$’s in (6.2)), but as soon as the conditions (6.2) are imposed, the initial velocities $\dot{x}_{i}(0)$ with respect to $t_{1}$ can not be arbitrary. Instead, the arbitrary parameters are the points $p_{i}$ which are singular points of the spectral curve. The integrals of motion of the Ruijsenaars-Schneider system for the $x_{i}$’s are determined by the spectral curve, so their values have to be expressed through the $p_{i}$’s. Therefore, we see that the problem should be posed in the following unusual way: given $x_{i}=x_{i}(0)$ and values of all higher integrals of motion, to find $\dot{x}_{i}(0)$ (which, according to (8.10), give $H_{i}$). This latter problem has more than one solution, and each solution corresponds to a common eigenstate of the quantum transfer matrices of the spin chain.
Taking all this into account, we can expect that eigenvalues of the Lax matrix (8.7) for the class of solutions constructed in Section 6 are expressed through the given parameters $p_{i}$ (which on the spin chains side are the twist parameters). To establish a precise relation, we use the differential-difference equation (5.15) for the adjoint wave function, i.e.,
$-\partial_{t_{1}}\psi^{*}(x+\eta,{\bf t};z)=\psi^{*}(x,{\bf t};z)+\partial_{t_{1}}\log\Bigl{(}\frac{\tau(x+\eta,{\bf t})}{\tau(x,{\bf t})}\Bigr{)}\,\psi^{*}(x+\eta,{\bf t};z),$ (8.11)
with $\psi^{*}$ given by (5.11) with the tau-function from (6.7). This function has $N$ simple poles at the points $x_{i}=x_{i}({\bf t})$. We can represent it as a sum of simple pole terms:
$\psi^{*}(x,{\bf t};z)=z^{-x/\eta}e^{-\xi({\bf t},z)}\left(c_{0}^{*}(z)+\sum_{i=1}^{N}\frac{c_{i}^{*}({\bf t},z)}{x-x_{i}}\right)$ (8.12)
with some coefficients $c_{i}^{*}$ parametrizing residues at the poles which can depend on $z$ and ${\bf t}$ but not on $x$. Plugging this ansatz into (8.11), we obtain the relation
$c_{0}^{*}+\sum_{i}\frac{c_{i}^{*}}{x-x_{i}+\eta}-z^{-1}\sum_{i}\frac{\dot{c}_{i}^{*}}{x-x_{i}+\eta}-z^{-1}\sum_{i}\frac{c_{i}^{*}\dot{x}_{i}}{(x-x_{i}+\eta)^{2}}$
$=c_{0}^{*}+\sum_{i}\frac{c_{i}^{*}}{x-x_{i}}+z^{-1}\sum_{j}\left(\frac{\dot{x}_{j}}{x-x_{j}}-\frac{\dot{x}_{j}}{x-x_{j}+\eta}\right)\left(c_{0}^{*}+\sum_{i}\frac{c_{i}^{*}}{x-x_{i}+\eta}\right)$
both sides of which are rational functions of $x$ with simple poles at $x=x_{i}$ and second order poles at $x=x_{i}-\eta$. It is easy to see that the highest order poles cancel identically and we need only to identify the residues at $x=x_{i}$ and $x=x_{i}-\eta$. For our purpose here it is enough to consider the poles at $x=x_{i}$. Their cancellation leads to the following system of linear equations for the coefficients $c_{i}^{*}$:
$zc_{i}^{*}+\dot{x}_{i}c_{0}^{*}+\dot{x}_{i}\sum_{j}\frac{c_{j}^{*}}{x_{i}-x_{j}+\eta}=0,\qquad i=1,\ldots,N.$ (8.13)
It can be written in a matrix form:
${\bf c}^{*}\dot{X}^{-1}\Bigl{(}zI-L\Bigr{)}=-c_{0}^{*}{\bf e},$ (8.14)
where ${\bf c}^{*}=(c_{1}^{*},\ldots,c_{N}^{*})$ is a row vector, $\dot{X}=\mbox{diag}\,(\dot{x}_{1},\ldots,\dot{x}_{N})$, $L$ is the Lax matrix (8.7) and ${\bf e}=(1,1,\ldots,1)$. The solution of this system is
${\bf c}^{*}({\bf t},z)=-c_{0}^{*}(z){\bf e}\Bigl{(}zI-L\Bigr{)}^{-1}\dot{X}.$ (8.15)
It remains to find $c_{0}^{*}(z)$. To this end, it is enough to tend $x\to\infty$ in (8.12) and (5.11) with the tau-function given by (6.7) and compare the results. This gives
$c_{0}^{*}(z)=z^{n}\prod_{i=1}^{n}(z-p_{i})^{-1}.$ (8.16)
By construction of the adjoint wave function (8.12) we know that it has multiple poles at $z=p_{i}$, $i=1,\ldots,n$. Therefore, looking at (8.15) with $c_{0}^{*}$ given by (8.16), we conclude that the eigenvalues of the Lax matrix should be identified with $p_{1},\ldots,p_{n}$, with each $p_{i}$ being in general multiple eigenvalue with multiplicity $M_{i}$ (then the right-hand side of (8.15) has a pole of order $M_{i}+1$ at this point, as it should).
Relating this result to quantum spin chains, we can reformulate it as follows. Consider the Lax matrix (8.7) $L(0)$ with the substitution $\dot{x}_{i}(0)=-\eta H_{i}$, where $H_{i}$ are eigenvalues (corresponding to a common eigenstate) of the quantum Hamiltonians ${\bf H}_{i}$ of the generalized twisted inhomogeneous spin chain given by (2.13):
$L_{ij}(0)=L_{ij}\Bigl{(}\{-\eta H_{l}\}_{N},\{x_{l}\}_{N}\Bigr{)}=\frac{\eta H_{i}}{x_{j}-x_{i}+\eta}.$
Then the spectrum of $L$ has the following specific form:
$\mbox{Spec}\,L\Bigl{(}\{-\eta H_{i}\}_{N},\{x_{i}\}_{N}\Bigr{)}=\Bigl{(}\underbrace{p_{1},\ldots,p_{1}}_{M_{1}},\,\underbrace{p_{2},\ldots,p_{2}}_{M_{2}},\,\ldots,\,\underbrace{p_{n},\ldots,p_{n}}_{M_{n}}\Bigr{)},$ (8.17)
where $M_{a}$ are eigenvalues of the operators ${\bf M}_{a}$ (2.10) on the eigenstate of the transfer matrix (we recall that ${\sum_{a=1}^{n}M_{a}=N}$) and $p_{i}$ are twist parameters (elements of the diagonal twist matrix ${\bf g}$). In other words, the characteristic polynomial of the Lax matrix is
$\det\Bigl{[}zI-L\Bigl{(}\{-\eta H_{i}\}_{N},\{x_{i}\}_{N}\Bigr{)}\Bigr{|}_{BE}\Bigr{]}=\prod_{i=1}^{n}(z-p_{i})^{M_{i}},$ (8.18)
where $L\Bigl{(}\{-\eta H_{i}\}_{N},\{x_{i}\}_{N}\Bigr{)}$ is taken on a solution to the Bethe equations. A very technical proof of this result, which essentially uses the nested Bethe equations (2.12), can be found in [16]. Here we have suggested another proof, which is easier and more instructive.
This result is also known as quantum-classical duality for integrable systems (quantum spin chain versus classical Ruijsenaars-Schneider). The remarkable relation between the two so different systems survives (and remains nontrivial) also in the limit $\eta\to 0$, where on the quantum side we have the Gaudin model and the Calogero-Moser system on the classical side (see [16, 17] for details).
From a more general viewpoint, the quantum-classical duality is a remarkable relation between the joint spectra of commuting quantum Hamiltonians and intersection of two Lagrangian submanifolds of the $2N$-dimensional phase space of an $N$-body classical integrable system of particles. A Lagrangian submanifold is an $N$-dimensional submanifold in the $2N$-dimensional phase space such that the restriction of the symplectic form $\omega=\displaystyle{\sum_{i=1}^{N}dP_{i}\wedge dx_{i}}$ to it is identically equal to zero. In the relation mentioned above, the first Lagrangian submanifold is the $N$-dimensional hyperplane corresponding to fixing all coordinates $x_{j}$ of the classical particles, while the second one is the level set of the $N$ independent integrals of motion in involution. Their dimensions are complimentary, and thus they intersect in a finite number of points. The essence of the quantum-classical duality is that the values of velocities $\dot{x}_{j}$ of the particles at the intersection points provide spectra of commuting quantum Hamiltonians of some quantum integrable model (one of the examples is the twisted inhomogeneous $GL(n)$-invariant spin chain considered in the present paper). Different intersection points correspond to different eigenstates of the commuting quantum Hamiltonians.
Let us say a few words about the meaning of this result. In particular, it makes it possible to solve the spectral problem for the quantum Hamiltonians without addressing the Bethe ansatz at any step. Instead, one should solve an “inverse spectral problem” for the Lax matrix of the classical integrable system of particles of Ruijsenaars-Schneider or Calogero-Moser type. Namely, let $\{x_{i}\}_{N}$ be inhomogeneity parameters of the spin chain and ${\bf g}=\mbox{diag}\,(p_{1},p_{2},\ldots,p_{n})$ its twist matrix. Let the eigenvalues of the Lax matrix be equal to the eigenvalues $p_{i}$ of the twist matrix, with some multiplicities $M_{i}$ such that ${\sum_{i=1}^{n}M_{i}=N}$. This fixes values of all the Ruijsenaars-Schneider integrals of motion according to (8.9). Then the spectrum of the non-local spin chain Hamiltonians ${\bf H}_{j}$ in the sector where eigenvalues of the weight operators ${\bf M}_{i}$ are equal to $M_{i}$ is given by the values of $H_{j}$ such that the matrix ${L_{ij}=\frac{\eta H_{i}}{x_{j}-x_{i}+\eta}}$ has the prescribed spectrum (8.17). This kind of duality suggests an alternative way to calculate joint spectra of commuting quantum transfer matrices without use of the coordinate or algebraic Bethe ansatz technique. There is also no need in such an unavoidable intermediate step as solving the Bethe equations. The spectra of quantum Hamiltonians appear to be encoded in algebraic properties of the Lax matrix for a very different purely classical model.
To be more precise, combining (8.5), (8.9) and (8.10), we obtain a system of algebraic equations for the joint spectrum of the Hamiltonians ${\bf H}_{i}$:
$\sum_{I\subset\{1,\ldots,N\},\,|I|=k}\,\,\Bigl{(}\prod_{i\in I}H_{i}\Bigr{)}\prod_{i<j,\,i,j\in I}\frac{(x_{i}-x_{j})^{2}}{(x_{i}-x_{j})^{2}-\eta^{2}}=e_{k}(\{\xi_{i}\}_{N}),$ (8.19)
where $e_{k}$ are the elementary symmetric polynomials. For example,
$e_{1}(\{\xi_{i}\}_{N})=\sum_{i}\xi_{i},\quad e_{2}(\{\xi_{i}\}_{N})=\sum_{i<j}\xi_{i}\xi_{j},\quad\mbox{and so on}.$
The set $\{\xi_{1},\ldots,\xi_{N}\}$ consists of $N$ given eigenvalues of the Lax matrix which are the twist parameters $p_{i}$ with multiplicities. In contrast to the Bethe equations, which are equations for some auxiliary variables, (8.19) are equations for the spectrum itself.
Finally, let us present the formulas which express the wave function, and its adjoint for polynomial solutions of the mKP hierarchy through the Lax matrix $L$ (8.7). We give them here without derivation (see [18, 44] for details):
$\begin{array}[]{l}\displaystyle{\psi(x,{\bf t};z)=\prod_{k=1}^{n}(1-p_{k}z^{-1})\,z^{x/\eta}e^{\xi({\bf t},z)}\frac{\det\Bigl{[}(xI-X)(zI-L)-\eta L\Bigr{]}}{\det\Bigl{(}xI-X\Bigr{)}\Bigl{(}zI-L\Bigr{)}},}\\ \\ \displaystyle{\psi^{*}(x,{\bf t};z)=\prod_{k=1}^{n}(1-p_{k}z^{-1})^{-1}z^{-x/\eta}e^{-\xi({\bf t},z)}\frac{\det\Bigl{[}(zI-L)(xI-X)+\eta L\Bigr{]}}{\det\Bigl{(}xI-X\Bigr{)}\Bigl{(}zI-L\Bigr{)}},}\end{array}$ (8.20)
where $X=X({\bf t})=\mbox{diag}(x_{1}({\bf t}),\ldots,x_{N}({\bf t}))$. Using (5.10), (5.11), one can see from (8.20) that the tau-function for this class of solutions is given by the following determinant formula:
$\tau(x,{\bf t})=\prod_{i=1}^{n}\Bigl{(}p_{i}^{x/\eta}e^{\xi({\bf t},p_{i})}\Bigr{)}\,\det_{N\times N}\left(xI-X_{0}+\eta\sum_{k\geq 1}kt_{k}L_{0}^{k}\right),$ (8.21)
where $X_{0}=X(0)$, $L_{0}=L(0)$.
9 Concluding remarks
We have reviewed the approach to quantum integrable models solvable by Bethe ansatz developed in [15][22], refining some arguments from these works and making the results more detailed. The essence of our approach is diagonalization of quantum transfer matrices by methods of the classical soliton theory, avoiding the Bethe ansatz procedure. It seems to us that the translation from the language of quantum integrability to the one of classical integrable hierarchies is suggestive and instructive and adds something important to the deeper understanding of both areas of mathematical physics. Objects and notions from the arsenal of the algebraic Bethe ansatz find their natural counterparts in the classical theory of the soliton equations. For example, factorization of the non-commutative generating function of transfer matrices for fundamental representations, which is a key step of the algebraic Bethe ansatz solution, on the classical side manifests itself as factorization of the difference wave operator of order $n$ into product of $n$ first order difference operators, each of them, being applied to the wave function, produces a Bäcklund transformation.
As is shown in [18], this approach, as well as the quantum-classical duality, can be extended to supersymmetric $GL(n|m)$-invariant spin chains (graded magnets) with rational $R$-matrices. However, the connection with the mKP hierarchy needs to be made more precise. Presumably, the wave operators relevant to $GL(n|m)$-invariant models are no longer finite difference operators but are of the form ${\bf W}(x)={\bf W}_{1}(x){\bf W}^{-1}_{2}(x)$, where ${\bf W}_{1}$, ${\bf W}_{2}$ are difference operators of orders $n$ and $m$.
In this paper we were restricted to quantum models with $R$-matrices that are rational functions of the spectral parameter. As we have seen, on the classical side they correspond to rational solutions of the integrable hierarchy, for which the tau-function is a polynomial or quasipolynomial. As it was argued in [20, 19], this approach can be extended, mutatis mutandis, to spin chains with trigonometric $R$-matrices (of the XXZ type), which are related to quantum deformations $U_{q}(gl_{n})$ of the universal enveloping algebras with a deformation parameter $q=e^{\gamma}$. (In these works it was assumed, however, that $q$ is not a root of unity.) On the classical side, such spin chains correspond to mKP tau-functions which are trigonometric polynomials of $x$ (i.e., Laurent polynomials of $e^{\gamma x}$). The solutions of this class, too, can be characterized by Krichever’s conditions. However, instead of (6.2) they have the form
$\sum_{m=-M_{i}/2}^{M_{i}/2}a_{im}\psi(x,{\bf t},p_{i}e^{2\gamma m})=0,$ (9.1)
where the sum goes over all integer numbers between $-M_{i}/2$ and $M_{i}/2$ for even $M_{i}$ and over all half-integer numbers between $-M_{i}/2$ and $M_{i}/2$ for odd $M_{i}$. The poles of the adjoint wave function of high orders $M_{i}+1$ at the points $p_{i}$ in the trigonometric case become “strings” of simple poles at the $M_{i}+1$ points $p_{i}e^{-2\gamma M_{i}},p_{i}e^{-2\gamma(M_{i}-1)},\ldots,p_{i}e^{2\gamma M_{i}}$. The case when $q$ is a root of unity is more complicated and requires a separate investigation. We hope to revisit the trigonometric case in a separate publication.
As far as a possibility to extend this approach to quantum models with elliptic $R$-matrices is concerned, it still remains to be a challenging open problem. On the first glance it seems that in this most general case the master $T$-operator can be defined by the same formula (4.1). However, one can see that the transfer matrices ${\sf T}_{\lambda}(x)$ with different $\lambda$’s have different monodromy properties under shifts by the periods. This means that the master $T$-operator defined by equation (4.1), being a linear combination of them, fails to be an “elliptic polynomial” of $x$, and so the existing theory of elliptic solutions to the mKP hierarchy is not applicable.
Finally, let us recall that the quantum-classical duality receives a formal “explanation” if one treats it as a limiting case of the so-called Matsuo-Cherednik correspondence, which was established in [51, 52] (see also [53, 54], where it was generalized and connected with the quantum-classical duality). The Matsuo-Cherednik correspondence, sometimes called the quantum-quantum duality, connects solutions to the (quantum) Knizhnik-Zamolodchikov equations (which can be regarded as a non-stationary extension of the spectral problem for the spin chain Hamiltonians ${\bf H}_{j}$) with stationary wave functions of the quantized Ruijsenaars-Schneider model. The parameter that controls non-stationarity in the former plays the role of the Planck’s constant in the latter. Tending it to zero makes the former problems stationary but still quantum, while the latter models become classical, and this is the way how the quantum-quantum (Matsuo-Cherednik) duality turns to the quantum-classical duality. As we have seen, the quantum-classical duality can be lifted to the level of integrable hierarchies of nonlinear equations like KP or mKP. It is then natural to ask whether something similar is true for the quantum-quantum duality, i.e., could it be obtained from a more general connection with an integrable hierarchy of KP or mKP type or its quantized version.
Acknowledgments
I am grateful to my co-authors A. Alexandrov, M. Beketov, A. Gorsky, V. Kazakov, I. Krichever, S. Leurent, A. Liashyk, O. Lipan, A. Sorin, Z. Tsuboi, P. Wiegmann, A. Zotov for collaboration in the works [2, 9], [15][19], to N. Reshetikhin and A. Veselov for discussions and to J. Harnad for careful reading of the manuscript and valuable remarks. Also, I would like to thank organizers of the Beijing Summer Workshop in Mathematics and Mathematical Physics at BIMSA, China (June 24 – July 5, 2024) for hospitality. This work has been supported in part within the state assignment of NRC “Kurchatov institute” (sections 2–4).